### v1

# Bridging AI and Science: Implications from a Large-Scale Literature Analysis of AI4Science

Abstract.
Artificial Intelligence has proven to be a transformative tool for advancing scientific research across a wide range of disciplines. However, a significant gap still exists between AI and scientific communities, limiting the full potential of AI methods in driving broad scientific discovery. Existing efforts in bridging this gap have often relied on qualitative examination of small samples of literature, offering a limited perspective on the broader AI4Science landscape. In this work, we present a large-scale analysis of the AI4Science literature, starting by using large language models to identify scientific problems and AI methods in publications from top science and AI venues. Leveraging this new dataset, we quantitatively highlight key disparities between AI methods and scientific problems in this integrated space, revealing substantial opportunities for deeper AI integration across scientific disciplines. Furthermore, we explore the potential and challenges of facilitating collaboration between AI and scientific communities through the lens of link prediction. Our findings and tools aim to promote more impactful interdisciplinary collaborations and accelerate scientific discovery through deeper and broader AI integration.
AI for Science, Large-Scale Literature Analysis, Link Prediction
1. Introduction
The 2024 Nobel Prizes in Physics and Chemistry were both awarded to Artificial Intelligence (AI) researchers. In particular, the developers of AlphaFold (Jumper et al., 2021) were recognized for their groundbreaking work, which is based on the Transformer network and has revolutionized the prediction of protein structures, providing a transformative approach to solving complex biological problems. This is just one example of how AI has emerged as a powerful tool to facilitate scientific research across various disciplines (Wang et al., 2023; Leontidis, 2024; Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
However, despite the transformative potential of using AI for advancing scientific research (AI4Science), a considerable gap persists between AI and scientific communities, hindering the full exploitation of AI for scientific discovery. On one side, advanced AI methodologies might remain mysterious or underutilized by scientists; on the other side, AI researchers may lack awareness of the specific challenges and potential applications in scientific domains, missing the opportunities for interdisciplinary collaborations.
Existing efforts to identify and bridge this gap have predominantly involved qualitative reviews of AI’s application in science (Wang et al., 2023; Leontidis, 2024), particularly in specialized areas like drug discovery and materials science (Mak et al., 2023; Guo et al., 2021; Xu et al., 2024). Based on qualitative examination of small samples of literature, these reviews often rely on heuristic insights from domain experts to suggest potential uses of AI in solving scientific problems. While valuable, such focused reviews are limited in providing a comprehensive and diverse perspective on the AI4Science landscape.
While recent research from the Science of Science community has conducted larger-scale studies that quantify AI’s impact on publications in scientific fields (Gao and Wang, 2023; Duede et al., 2024), they often heavily rely on established scientific taxonomies in “AI-heavy” fields. There remains a lack of delivering a holistic, dynamic, and data-driven overview of AI4Science. Such analysis is crucial for understanding the barriers and identifying new opportunities for deep engagement of fast-evolving AI methodology in scientific research.
In this work, we aim to depict the AI4Science landscape via a large-scale and comprehensive analysis of relevant literature in both scientific and AI communities over the past decade. We start by using large language models (LLMs) to identify the scientific problems and AI methods addressed in publications from top science and AI venues, through which we assemble a novel and balanced AI4Science dataset to analyze the role of AI in scientific research (Sec. 3). Then with projection maps (Fig. 1) and a bipartite graph (Fig. 2) constructed from this dataset, we quantitatively reveal the discrepancies between AI and science, leading to several key findings and implications (Sec. 4). Finally, through the lens of link prediction, we demonstrate the potential and challenges in advancing the bridge between AI and science (Sec. 5). We anticipate that our findings and tools will foster more effective interdisciplinary collaborations, ultimately accelerating scientific discovery through a more diverse and deeper integration of AI methods.
Figure 1. Illustration of the AI4Science dataset: (1) LLM-based extraction of scientific problem $p_{i}$, AI method $m_{i}$, and AI usage $u_{i}$ from an example publication. (2) 2D projections of the extracted scientific problems $\{p_{i}\}$ and AI methods $\{m_{i}\}$. Each dot represents a publication, with coordinates obtained by reducing the dimensionality of the problem/method embeddings. (a) Green dots correspond to AI4Science work, where AI is utilized to address scientific problems; Orange dots are papers addressing scientific problems with non-AI techniques; (b) Purple dots indicate using AI to address non-scientific problems. The visualizations reveal a noticeable discrepancy in the distribution of AI4Science work (green) versus non-AI4Science work (orange and purple) in both the problem and method spaces.
Figure 2. The AI-Science bipartite graph and node distributions. (a) The bipartite graph with scientific problem clusters and AI method clusters as nodes. The size of each node corresponds to its unweighted degree, representing the number of AI methods applied to a scientific problem, or the number of scientific problems that applied an AI method. For clarity, we omit edges representing fewer than five publications, and isolated nodes are removed thereafter. The distribution of links indicates an imbalanced connectivity. (b-c) The degree of AI method nodes follows a log-normal distribution, and the degree distribution of scientific problem nodes is even more heavily tailed, suggesting the existence of “hubs” in linking AI and science.
Notation Value
Publication year range 2014–2024
# of publications $N$ 159,295
# of scientific problem extractions |$|\{p_{i}\}|$ 123,773
# of scientific problem clusters |$|\{P_{i}\}|$ 369
# of AI method extractions |$|\{m_{i}\}|$ 35,522
# of AI method clusters |$|\{M_{i}\}|$ 345
# of AI usage extractions |$|\{u_{i}\}|$ 35,522
# of AI4Science publications 7,024
Table 1. AI4Science dataset $\mathcal{D}=\{(p_{i},P_{i},m_{i},M_{i},u_{i})\}_{i=1}^{N}$.
2. Related Work
2.1. Qualitative Literature Reviews of AI4Science
To identify potential applications of AI techniques in science, researchers typically turn to qualitative literature reviews of relevant papers. These reviews include general explorations of AI’s usage across various scientific domains (Wang et al., 2023; Leontidis, 2024), as well as more focused examinations in specific fields such as drug discovery (Dara et al., 2022; Mak et al., 2023; Blanco-Gonzalez et al., 2023; Qureshi et al., 2023), materials design (Sha et al., 2020; Li et al., 2020; Guo et al., 2021), and social sciences (Xu et al., 2024).
While these reviews provide valuable insights, they often rely on interdisciplinary experts to manually identify and categorize scientific problems and potential AI methods, limiting their scope. Such an approach often could not offer a comprehensive and data-driven overview of the broader “AI4Science” landscape.
2.2. Quantitative Literature Analysis
In addition to heuristic and qualitative literature reviews, a few recent papers from the Science of Science community have focused on quantitatively analyzing the AI4Science publications. For instance, Gao and Wang (2023) examined how AI benefits scientific research, finding that AI use is prevalent across various scientific disciplines, and that papers incorporating AI tend to receive a citation advantage. Similarly, Duede et al. (2024) explored AI’s engagement in different fields, uncovering an “oil-and-water” phenomenon – while AI-engaged work is spreading across disciplines, it does not integrate well with non-AI-engaged work within each specific field.
Although these studies offer valuable insights into AI4Science, they primarily focus on areas where AI and science are already “blended” effectively, and overlook the broader and underexplored parts of the landscape. As a result, these approaches often fail to offer insights into the disparities between AI and scientific communities. In contrast, our work offers a more comprehensive view of the AI4Science landscape by showing and quantifying the connections and gaps between scientific problems and AI methods, as well as insights to uncover potential collaborations.
3. A New Dataset of AI4Science
A major challenge in understanding the disparities and potential of AI4Science is the lack of a comprehensive, balanced dataset. Existing literature data are often biased toward success stories, such as areas where AI and science are already deeply intertwined. Or they reflect the perspective of either the scientific or AI communities, but not both. For example, the AI methods discussed in the scientific literature may not represent the full range of advancements in AI, while the scientific applications explored in AI literature may fail to capture the nuances and challenges of real-world scientific problems. We take the initiative to create a broad and balanced literature dataset to provide a wide-angled lens for understanding AI4Science research.
3.1. Curation of AI4Science Literature
We collect publications from leading science and AI venues to offer a balanced view from both communities:
  • For science domains, we include three top multidisciplinary journals that represent cutting-edge and high-quality research (Nature, Science, and PNAS) and two of their subjournals (Nature Communications, and Science Advances).
  • For AI communities, we include six top conferences from the list of CSRankings.org, comprising two AI-focused venues (AAAI, IJCAI), three machine learning venues (ICLR, ICML, NeurIPS), and the Web Conference (WWW) as a venue of applied AI research.
Our data spans the past decade, from 2014 to 2024, including a total of 159,295 publications from both AI and science venues. More details of the data curation process can be found in Appendix A.1.
3.2. Extracting Scientific Problems and AI Methods with LLMs
To understand how AI is applied to science, we need to extract paired entries of scientific problems and AI solutions from relevant publications. Traditionally, this is accomplished through literature reviews and surveys which involve manual examinations of small samples of papers (Wang et al., 2023; Leontidis, 2024), a labor-intensive and time-consuming process. Recent quantitative studies in the Science of Science field have started using metadata entries, such as keywords, to identify AI usage and categorize scientific topics (Gao and Wang, 2023; Duede et al., 2024). While this method enables larger-scale analysis, it offers only a coarse-grained view, potentially missing the nuanced details of scientific problems and AI methods. Additionally, it can suffer from issues such as incomplete data entries, limiting the effectiveness of analysis.
In contrast, our work leverages large language models (LLMs) to extract detailed descriptions of scientific problems and AI methods from publication titles and abstracts, allowing for a more nuanced, scalable, and data-driven analysis. Such LLM-based extraction techniques have also been employed in recent studies for extracting aspects from scientific literature, such as Zhang et al. (2024).
Particularly, for each publication of interest, we use OpenAI GPT-4o mini to extract the following three key aspects:
  • Scientific Problem ($p_{i}$): The primary scientific problem addressed in the paper, including both a keyword or keyphrase summarizing the problem and a detailed definition.
  • AI Method ($m_{i}$): The main AI method applied in the paper, including a keyword or keyphrase summarizing the method and a detailed description.
  • AI Usage ($u_{i}$): A detailed explanation of how the AI method is specifically applied to the scientific problem.
It is possible that a publication did not address a scientific problem or did not use an AI method, in which case the corresponding field(s) would remain empty. If a paper addresses a scientific problem using an AI method – i.e., $p_{i}$, $m_{i}$, and $u_{i}$ are all non-empty – it is considered as an AI4Science work. Figure 1 presents an example of the extraction results of an AI4Science publication.
To assess the reliability of GPT extractions, we conducted a small-scale human evaluation on 100 papers. Each extraction record was reviewed by at least two annotators, resulting in an average accuracy of 91.0%. For more details on the extraction prompts, additional examples of the extracted data, and verification with human annotations, please refer to Appendix A.2.
3.3. Semantic Clustering of Scientific Problems and AI methods
One of the key differences between this new AI4Science dataset and the data used in existing studies (Gao and Wang, 2023; Duede et al., 2024) is that our dataset contains detailed textual descriptions of the identified scientific problems and AI methods rather than pre-defined taxonomies, allowing for a deeper semantic analysis of their relationships.
To achieve this, we generate semantic embeddings for the scientific problems ${p_{i}}$, AI methods ${m_{i}}$, and AI usages ${u_{i}}$ using the InstructorEmbedding model (Su et al., 2022). We then apply LargeVis (Tang et al., 2016) to project these high-dimensional embeddings into 2D coordinates. Next, HDBSCAN (Campello et al., 2013) is utilized to group similar scientific problems and AI methods. For each cluster, we use GPT-4o to summarize the main topic based on frequent keywords and samples of extraction results. This process results in 369 and 345 clusters respectively, with corresponding cluster labels $P_{i}$ and $M_{i}$ to each data point. Fig. 1 displays the 2D visualizations, and Fig. 4-5 in the Appendix presents the maps with cluster names annotated. Details of the process can be found in Appendix A.3.
Unlike directly using keywords or categorizations from publication metadata (Gao and Wang, 2023; Duede et al., 2024), our approach employs a data-driven methodology. By using clustering and summarization, relevant AI methods discussed in both scientific and AI literature are grouped together, as are the scientific problems mentioned in both fields. This ensures greater consistency and generalizability of the data entries.
3.4. Dataset Overview and the Bipartite Graph
Through LLM-based extraction and semantic clustering, we form a comprehensive dataset, $\mathcal{D}=\{(p_{i},P_{i},m_{i},M_{i},u_{i})\}_{i=1}^{N}$, which includes the scientific problems, AI methods, their corresponding cluster labels, and the usage of AI for science. The basic statistics of this dataset are listed in Table 1.
To better illustrate the connections between AI and science, we construct a bipartite graph based on the clusters. In this bipartite graph, scientific problem clusters and AI method clusters represent two types of nodes, and the publications act as edges connecting them. Formally, the graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{P_{i}\}\cup\{M_{i}\}$ and $\mathcal{E}=\mathcal{D}$. As shown in Fig. 2(a), the graph provides an intuitive visual representation of how AI is linked to scientific challenges and vice versa.
(a) Scientific problem clusters.
(b) AI method clusters.
Figure 3. Scientific problem clusters (a) and AI method clusters (b) are visualized as circles in the scatter plots. In both plots, the $x$-axis represents the total number of publications in each cluster, while the $y$-axis reflects the number of interdisciplinary AI4Science publications in each cluster. Three outstanding points are skipped for clarity1. The black lines show the regression results on the clusters. Clusters falling below the regression line highlight the under-investigated regions, where the integration of AI and science remains limited. These areas represent potential opportunities for further interdisciplinary collaboration.
4. Discrepancy Between AI and Science
Both Fig. 1 and 2 show a noticeable pattern of discrepancy between AI and science. This section quantitatively assesses these disparities, providing a deeper understanding of the gaps.
4.1. Uneven Distribution of AI4Science Research
The projection maps in Fig. 1 illustrate the distribution of AI4Science work within the semantic spaces of scientific problems and AI methods (the green dots among orange and purple dots). The uneven distributions suggest that AI4Science work is heavily clustered in certain subareas, while sit remains underrepresented in wide ranges of both the problem and the method spaces.
In Fig. 3, we plot the clusters of scientific problems and AI methods, showing the relation between cluster sizes and the presence of AI4Science publications within each cluster. The slopes of the regression lines represent the average proportions of AI4Science work in the literature of each scientific problem cluster or AI method cluster. Clusters below the regression line indicate problems or AI methods in which the integration of AI4Science is underrepresented (noted the under-investigated regions), suggesting potential opportunities for further interdisciplinary collaboration.
Under-investigated scientific problems.
In the bottom-right section of Fig. 3(a), below the regression line, several scientific problem clusters are highlighted as being rarely addressed by AI techniques. Notable examples include Magnetic Materials, Asymmetric Synthesis, Lithium-Ion Battery Challenges, Antibiotic Resistance, Stem Cell Differentiation, Viral Mechanisms, Climate Change Emissions, and Enzyme Catalysis. These scientific problems could benefit from more engagement of AI methods.
Under-investigated AI methods.
Similarly, Fig. 3(b) reveals AI techniques that have not yet been widely applied to sciences, including: Attention Mechanism, Gradient-Based Methods/Gradient Descent, Embedding Techniques, Policy Optimization, Regularization Techniques, Theoretical Analysis, Adversarial Training, Contrastive Learning, and Variational Inference. These techniques could have been explored in a wider range of scientific applications.
11footnotetext: Three outstanding data points are skipped from the scatter plots for visualization clarity: Neural Network Challenges (838, 653) in (a), Machine Learning (601, 446) and Deep Learning (577, 308) in (b).
Node type Partition Avg. Deg.
Scientific problem clusters (n=369) All nodes 15.7
Well-investigated 63.2
Under-investigated 6.3
AI method clusters (n=345) All nodes 16.5
Well-investigated 29.2
Under-investigated 8.5
Table 2. Average degree of nodes in the bipartite graph.
4.2. Uneven Distribution of Bipartite Links
Fig. 2 reveals an uneven distribution of node sizes in the bipartite graph, indicating the presence of “hub” and “peripheral” nodes among both scientific problems and AI methods. This implies that certain scientific challenges and AI techniques play a central role in interdisciplinary AI4Science work, while others are less explored.
Quantitatively, heavy-tailed degree distributions are observed as visualized in Fig. 2(b-c). Specifically, the degrees for AI method nodes follow a log-normal distribution, while the degree distribution for scientific problem nodes exhibits an even heavier tail, though not as extreme as a power-law distribution. These findings support the idea that a small number of “hubs” are linked to a large variety of AI methods or scientific problems, while many other nodes remain peripheral to these interdisciplinary connections. Heavy-tailed distributions, particularly log-normals, are also commonly observed in other bipartite networks (Vasques Filho and O’Neale, 2018), such as document-term graphs (Wang et al., 2014), and in collaborative networks (Perc, 2010; Sun et al., 2014). This suggests that the structure of the AI4Science bipartite graph reflects a general pattern of interdisciplinary and collaborative work.
Table 12 in the Appendix lists the “hubs” with the highest node degrees, representing the most interconnected scientific problems and AI methods. From the table, we observe that the scientific problems linked to the widest variety of AI methods are primarily computational or data-processing problems, such as: Neural Network Challenges, Statistical Inference, Language and Auditory Processing. These are followed by more domain-specific problems like: Social Media Dynamics, Decision-Making Processes, Market Economics. On the AI methods side, the techniques applied to the greatest number of scientific challenges include general methods such as: Machine Learning, Deep learning, Neural Networks, Classification Methods, Data Analysis, Bayesian Inference. In addition, there are specialized techniques tailored to scientific domains, including: Genomic Analysis, Computational Biology, Protein Design, Molecular Modeling, Material Informatics.
We further break down the node degree distribution based on cluster partitions. Specifically, we calculate the average node degree for well-investigated and under-investigated clusters, as identified earlier in Sec. 4.1. The statistics are listed in Table 2. We observe a higher average degree for well-investigated scientific problems and AI methods. This suggests that under-investigated regions not only have fewer AI4Science publications, but also exhibit a lesser variety, or limited connectivity in the bipartite graph.
4.3. Discrepancy Between the AI and Science Communities
Our dataset includes publications from both scientific journals and AI conferences, allowing us to explore how these two communities approach the integration of AI into science in distinct ways.
Table 13 lists the top clusters emphasized by each community. In terms of facilitating scientific discovery with AI, the science community places greater emphasis on challenges such as: Protein Structure and Design, Materials Design, Artificial Intelligence Ethics, and Single-Cell RNA Sequencing; In contrast, the AI community tends to focus more on areas like: Social Media Dynamics, Game Theory, Market Economics, etc. When it comes to applying AI methods to scientific problems, the two communities also show different preferences. In addition to general approaches like Machine Learning and Deep Learning, the science community frequently employs AI techniques specifically tailored for scientific challenges, such as: Genomic Analysis, Protein Design, and Computational Biology. In comparison, the AI community utilizes a broader range of general AI methodologies, covering diverse areas such as: Data Analysis, Game Theory, Causal Inference, Probabilistic Modeling, Reinforcement Learning, Optimization Methods, Generative Models, etc.
In addition, by comparing Table 13 with Table 12, we also find that the scientific problems addressed by the AI community, tend to be connected to a wider variety of AI methods. Similarly, AI methods frequently used by the science community are applied across a broader range of scientific challenges. Similarly, AI methods frequently used by the science community are also more likely to be applied across a broader range of scientific challenges. This highlights the discrepancy between the two communities, where the science community engages with a limited range of AI methods compared to the AI community, and vice versa.
The distribution of AI4Science work published in scientific journals and AI conferences in the semantic maps (Fig. 6 in the Appendix) also illustrates the different focuses of these two communities.
4.4. Findings and Implications
Through our analysis of the distribution of AI4Science work, bipartite graph node degrees, and the discrepancies between the AI and science communities, we derive the following key findings:
  1. (F1)
    Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods under-investigated in the collaboration context (Sec. 4.1);
  2. (F2)
    The connectivity between scientific problems and AI methods is highly imbalanced, with certain nodes acting as “hubs” while other peripheral nodes are less connected (Sec. 4.2);
  3. (F3)
    The science and AI communities take distinct approaches to integrate AI into scientific research, prioritizing different problems and methods (Sec. 4.3).
These key findings lead to two important implications for fostering a wider and deeper exploration of AI4Science:
  1. (I1)
    More attention should be directed toward exploring the under-investigated areas, incorporating a broader range of scientific challenges and AI techniques into the AI4Science landscape.
  2. (I2)
    Efforts should be made to discover new connections between AI and science, uncovering innovative ways to apply AI methods to scientific research.
5. Bridging AI and Science: Through the Lens of Link Prediction
Building on the findings and implications above, we delve deeper into the potentials and challenges of advancing the connections between AI and science from the perspective of link prediction.
Setting Method Precision Recall F1
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Sci $\to$ AI Katz index 0.430 0.335 0.320 0.272 0.093 0.223 0.311 0.450 0.153 0.267 0.315 0.339
Node2vec 0.988 0.836 0.719 0.530 0.259 0.505 0.625 0.765 0.411 0.630 0.669 0.626
LLM+RAG 0.326 0.216 0.200 0.140 0.058 0.096 0.129 0.150 0.099 0.132 0.157 0.145
LLM+Graph 0.541 0.594 0.530 0.437 0.137 0.311 0.419 0.558 0.218 0.408 0.468 0.490
AI $\to$ Sci Katz index 0.485 0.402 0.362 0.291 0.095 0.205 0.262 0.344 0.160 0.271 0.304 0.316
Node2vec 0.997 0.904 0.837 0.717 0.197 0.371 0.470 0.622 0.329 0.526 0.602 0.666
LLM+RAG 0.368 0.293 0.267 0.210 0.036 0.057 0.077 0.098 0.066 0.096 0.120 0.133
LLM+Graph 0.458 0.316 0.300 0.201 0.068 0.113 0.163 0.152 0.118 0.167 0.211 0.173
Table 3. Link prediction results of different models.
Method Link prediction setting Precision Recall F1
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Node2vec Sci $\to$ AI Well-investigated 1.000 0.980 0.922 0.799 0.115 0.295 0.417 0.606 0.207 0.453 0.574 0.689
Under-investigated 0.960 0.726 0.535 0.310 0.392 0.734 0.841 0.928 0.557 0.730 0.654 0.465
AI $\to$ Sci Well-investigated 0.996 0.956 0.916 0.807 0.126 0.293 0.414 0.616 0.224 0.448 0.571 0.699
Under-investigated 0.978 0.651 0.481 0.284 0.560 0.836 0.918 0.978 0.712 0.732 0.631 0.440
LLM+RAG Sci $\to$ AI Well-investigated 0.422 0.271 0.247 0.189 0.031 0.054 0.074 0.102 0.057 0.090 0.113 0.132
Under-investigated 0.240 0.178 0.157 0.111 0.081 0.149 0.182 0.224 0.121 0.162 0.169 0.148
AI $\to$ Sci Well-investigated 0.344 0.259 0.267 0.206 0.027 0.053 0.076 0.098 0.050 0.088 0.118 0.133
Under-investigated 0.188 0.115 0.089 0.066 0.090 0.140 0.159 0.167 0.122 0.126 0.114 0.095
Table 4. Link prediction results on well- and under-explored regions.
5.1. Methodology
In the curated AI4Science dataset $\mathcal{D}=\{(p_{i},P_{i},m_{i},M_{i},u_{i})\}_{i=1}^{N}$, the extracted and clustered scientific problems $\{p_{i}\}$ and $\{P_{i}\}$, as well as AI methods $\{m_{i}\}$ and $\{M_{i}\}$, represent the respective landscapes of these two domains. The AI4Science publications, along with their descriptions of AI usage $\{u_{i}\}$, highlight the connections between scientific challenges and AI methods. To model these connections, we employ the link prediction formulation.
Data.
We split the data into training and test sets based on publication dates. The training set includes papers published between 2014 and 2022 (5,829 AI4Science publications), which are used for training models, retrieving data points for generation augmentation, and identifying well/under-investigated areas, as illustrated in Fig. 3. The testing set consists of publications from 2023 and 2024 (1,195 AI4Science publications), serving as the ground truth for evaluating link prediction models. The statistics are provided in Table 14 in the Appendix. It is important to note that the analyses prior to this section involve the full AI4Science dataset, spanning the period from 2014 to 2024.
Model and evaluation.
We employ two types of models to predict links between scientific problems and AI methods. The first category includes conventional bipartite link prediction models (Özer et al., 2024), such as those based on the Katz index (Katz, 1953) and node2vec embeddings (Grover and Leskovec, 2016). Formally, given a source node $P_{i}$ (or $M_{i}$) in the bipartite graph, we predict $K$ potential target nodes for linking. We then compare the predicted set with the ground truth target node set and report Precision, Recall, and F1 scores @K (Al Hasan et al., 2006; Yang et al., 2015). The bipartite graph is constructed and the models are trained using only the training data, with evaluation performed on the testing data. Please note that existing links may also be counted in constructing the ground truth set, because a new AI4Science work may reuse an existing connection between a scientific problem and an AI method.
In addition to conventional link prediction, we utilize a large language model (LLM) for generative predictions. Specifically, given a scientific problem $p_{i}$ from cluster $P_{i}$, we generate potential AI methods $m_{i}$ that could be applied. These generated texts are then mapped to cluster labels $M_{i}$ through embedding similarity matching. Similarly, given an AI method $m_{i}$, we can generate potential scientific problems $p_{i}$ that it could address. We use the OpenAI gpt-3.5-turbo-0125 model as the generator. Since this model’s knowledge is current up to September 2021, there is no risk of data leakage. The prediction process is further enhanced by retrieving similar papers from the training set using retrieval-augmented generation (RAG) (Lewis et al., 2020).
While we can use large language models for generative predictions leveraging embedding similarity matching, the large language model is only exposed to a small neighborhood of literature through RAG and has no access to the higher-level linked bipartite graph structure. Thus, we include results where the entire graph (training links) is provided to the model. Given a scientific problem cluster $P_{i}$, all possible AI method clusters $M$ and past links between the clusters, we directly generate AI method cluster labels $M_{i}$. Similarly, given a AI methods cluster $M_{i}$, all possible scientific problem clusters $P$ and past links between the clusters, we directly generate scientific problem cluster labels $P_{i}$.
For more detailed descriptions of the models and evaluation metrics, please refer to Appendix C.
5.2. Overall Link Prediction Results
Table 3 presents models’ performances on the test data, where conventional link prediction methods show strong results in both Precision and Recall. Particularly, while both Katz and node2vec rely on suggesting similar nodes to link, node2vec outperforms Katz, suggesting the underlying mechanism connecting scientific problems and AI methods goes beyond homophily in the local neighborhood and potentially considers the broader structures of the AI4Science network.
Notably, the LLM+RAG approach lags behind conventional link prediction methods, emphasizing the challenge of generative link prediction at the individual paper level. A potential explanation is that, unlike node-level link prediction in a bipartite graph, LLMs struggle to obtain a comprehensive global view of the AI4Science landscape, even with the aid of RAG. This implies the benefit of considering the global bipartite graph structures and the opportunity to build knowledge graphs in the AI4Science domain to help LLMs better understand the broader context.
This is further elaborated by the results for LLM+Graph, where the model is provided with global bipartite graph structure. LLM+Graph outperforms LLM+RAG, suggesting that LLM benefits from bipartite linked graph information. Moreover, LLM+Graph achieves comparable results with Katz index, which implies LLM’s ability to understand graph-level features and make informed link predictions.
Nevertheless, LLM+RAG offers a unique advantage: it can generate detailed ideas in texts for AI4Science applications. Table 19 in the appendix reports evaluation results on text generation metrics, where LLM+RAG demonstrates notable results.
5.3. Prediction in the Well- and Under-investigated Regions
Fig. 3 reveals the presence of well- and under-investigated areas in both the scientific problem and AI method spaces. Echoing Implication (I1), we comparatively explore the potentials and challenges in facilitating interdisciplinary work for these regions.
Formally, we refer to scientific problem clusters or AI method clusters above the regression lines as well-investigated areas, while those that fall below the lines are the under-investigated regions. Clusters that reside within the 95% confidence intervals of the regression results are excluded from this analysis.
Table 4 compares the link prediction results for well-investigated and under-investigated clusters. In both categories of link prediction methods, we observe a similar trend: Precision scores are generally higher for well-investigated scientific problems and AI methods, even as K increases. In contrast, Recall is relatively higher for under-investigated clusters. These results are consistent with the node degree statistics in Table 2, where the average degree of under-investigated nodes is much lower, making it easier for the models to predict links, resulting in higher Recall. For instance, node2vec can retrieve nearly all ground truth links with K=10.
Additionally, in the under-investigated partition, the rapid decline in Precision as K increases suggests that link prediction models are capable of discovering novel links that have not been identified by human researchers in the test data. This highlights great opportunities for expanding the connections between AI methods and scientific challenges.
Reference Method # of novel (unseen) links
@1 @3 @5 @10
Train Node2vec 254 820 1,369 2,722
LLM+RAG 870 2,502 3,983 6,775
Train+Test Node2vec 4 396 899 2,203
LLM+RAG 760 2,308 3,748 6,505
Table 5. Number of novel links discovered by models compared with the training set and the whole dataset. The test data introduced 554 new links compared to the training data.
5.4. Discovering Novel Links
Previous discussions highlight the capability of link prediction models to uncover novel links. Building on this, we further investigate the newly discovered connections in response to Implication (I2).
Compared to the training data up to 2022, human researchers identified 554 new links between scientific problems and AI methods during 2023 and 2024. In contrast, link prediction methods have demonstrated the potential to discover even more connections, especially as K increases, as shown in Table 5.
Remarkably, the LLM+RAG approach demonstrates strong potential in generating new links, even with a relatively small number of predictions per data point (e.g., @1). This observation aligns with recent research suggesting that LLMs are capable of proposing novel research ideas (Si et al., 2024). Many of the predicted links involve applying Machine Learning, Deep Learning, and various types of Neural Networks (e.g., GNN, RNN, GAN), as well as Reinforcement Learning, to address a wide range of scientific challenges.
In comparison, conventional link prediction methods like node2vec, which leverage cluster-level bipartite graph structures, exhibit high accuracy in identifying new links in the test data but show lower novelty compared to LLM+RAG. For instance, in the @1 predictions, the node2vec model successfully identifies 250 of the 254 new links (98.4%) that were also discovered by human researchers. The four remaining links, not identified by human researchers, include: Cardiac Pathophysiology – Matrix Factorization, Mycobacterium Tuberculosis – Contrastive Learning, Plant Stress Responses – Topological Data Analysis, and Prostate Cancer – Ensemble Learning. Some of these links have since been developed into publications (Wang et al., 2019; Tse et al., 2020; Palande et al., 2023; Patel et al., 2024), highlighting the model’s ability to suggest valuable connections ahead of human exploration.
6. Limitations, Challenges, and Future Work
Potential bias in data selection.
Our analysis focuses on publications from leading science and AI venues, specifically top-tier journals and conferences. While this ensures high-quality and influential work, it introduces a selection bias that may overlook significant contributions from less prominent venues, potentially resulting in an incomplete representation of the AI4Science landscape. To mitigate this bias, future work could expand the dataset to include a broader range of publications. Given that our analysis framework is flexible and can easily accommodate new data sources, such an expansion would be a valuable direction for achieving a more comprehensive understanding of the AI4Science field.
Abstract-based extraction.
The extraction of scientific problems, AI methods, and AI usages in this study rely solely on paper titles and abstracts, due to limited access to full texts and cost considerations. This approach aligns with the existing AI4Science literature analysis efforts, which also focus on the title and abstract as in the metadata (Gao and Wang, 2023; Duede et al., 2024). While this method provides useful insights, it may miss important details present in the full text of publications. For example, when extracting the AI methods used in the AlphaFold paper (Jumper et al., 2021), as illustrated in Fig. 1, GPT identifies only the general use of a Neural Networks approach from the abstract. However, the specific use of the backbone Transformer model is detailed in the full text. In future work, efforts can be directed toward incorporating full-text analysis to enrich the extracted content and capture more nuanced AI methodologies used in scientific research.
The definition of “AI4Science”.
There is no universally accepted standard for defining the boundaries of the broad terms “AI”, “science”, and “AI4Science”, which presents a challenge for accurate classification. Previous studies have relied on keywords and human annotations (Gao and Wang, 2023; Duede et al., 2024), but this method can overlook less-known techniques and lacks the ability to capture the full semantic space of AI methods used in science. In contrast, our approach leverages the knowledge and capabilities of LLMs to classify AI4Science work, with human verification on a small-scale subset. To foster more consistent and scalable analyses, the AI4Science community should work toward developing standardized classification guidelines for automatic large-scale literature analysis.
Evaluation of link prediction.
Another limitation of our study is the reliance on publication data as the ground truth for evaluating link prediction models. While this provides a reliable benchmark, it may not fully capture the breadth of potential AI4Science connections, and overlook novel or unconventional links that the models might suggest. Future work could consider incorporating more comprehensive evaluation strategies, such as expert reviews or real-world validation, to assess the effectiveness of the link predictions.
Limited exploration in link prediction models.
Our primary goal in this study is not to optimize the link prediction model, but rather to use link prediction as a lens to explore the potential of connecting AI methods and scientific problems. As a result, we only included a few link prediction models, leaving room for further research to improve the performance of AI4Science link prediction or building recommender systems for AI methods or scientific problems.
7. Conclusion
This paper aims to identify and bridge the gaps between AI and science in the context of AI4Science research. We introduce a comprehensive, large-scale dataset of AI4Science publications, with scientific problems and AI methods extracted using large language models (LLMs). Through quantitative analysis of this dataset, we uncover several key disparities: (1) Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods under-investigated; (2) The connectivity between scientific problems and AI methods is highly imbalanced, with certain nodes acting as “hubs” while other peripheral nodes are less connected; (3) The science and AI communities take distinct approaches to integrating AI into scientific research, prioritizing different problems and methods. We further investigate the potential and challenges of fostering AI4Science collaboration through the lens of link prediction. The experiment results demonstrate great opportunities for using link prediction models to explore the under-investigated scientific problems and AI methods, as well as discover novel connections. We anticipate that our findings and tools will provide valuable insights to enhance interdisciplinary collaboration and accelerate scientific discovery through the integration of AI.
Acknowledgements.
The authors would like to extend their sincere appreciation to Shixuan Liu and Sicheng Shen for their contributions as annotators.
Appendix A AI4Science Dataset
A.1. Curation of Literature
Data source.
  • For the science domain, we obtained our data via the NCBI Entrez database, encompassing the following papers: Nature, Science, PNAS, Nature Communications, Science Advances. Only papers with both title and abstract available and within 2014 to 2024 are kept. A brief statistics of science domain papers are as follows:
    Journal #Pub. #AI4Science Pub.
    Nature 10,890 254
    Science 10,998 207
    PNAS 35,556 997
    Nature Communications 54,243 2,027
    Science Advances 12,086 488
    Total 123,773 3,973
    Table 6. Science Domain dataset statistics
  • For AI domain, we obtained our data from MASSW (Zhang et al., 2024), which encompasses the following six top AI conferences: AAAI, IJCAI, ICLR, ICML, NeurIPS, WWW. Only papers with both title and abstract available and within 2014 to 2024 are kept. A brief statistics of AI domain papers are as follows:
    Conference #Pub. #AI4Science Pub.
    NeurIPS 11,250 889
    AAAI 8,994 915
    ICML 5,641 530
    WWW 3,777 241
    IJCAI 2,953 271
    ICLR 2,907 205
    Total 35,522 3,051
    Table 7. AI Domain dataset statistics
A.2. Extraction with LLMs
Extracting key aspects
As aforementioned, we asked large language models to extract the following three key aspects: Scientific Problem, AI Method, Usage. To help LLM better understand the three aspects, we further break down the three key aspects into the following six aspects when prompting the model:
  • Scientific Problem (keyword/keyphrase): A few keywords or key phrases that summarize the primary scientific problem addressed in the paper.
  • Scientific Problem (definition): A detailed definition of the primary scientific problem addressed in the paper.
  • Scientific Problem discipline: The scientific discipline that the primary scientific problem addressed in the paper falls into.
  • AI method (keywords/keyphrase): A few keywords or key phrases that summarize the AI method used in the paper.
  • AI method (definition): A detailed definition of the AI method used in the paper.
  • AI Usage: A detailed explanation of how the AI method is specifically applied to the scientific problem.
Prompting Template
We used the following prompt to extract the six key aspects of a paper:
## Background and Task Description
You are an expert in both science and artificial intelligence (AI), where AI generally refers to intelligence exhibited by machines (particularly computer systems), including models, algorithms, etc. Given a scientific paper, your task is to extract the following aspects from the title and abstract:
**Problem (keyword/keyphrase):** A keyword or a keyphrase that summarizes the main problem to be addressed in this paper.
**Problem (definition):** The detailed definition of the problem.
**Problem discipline:** The discipline in which the main problem best fits.
**Method (keyword/keyphrase):** A keyword or a keyphrase that summarizes the main method used in this paper to address the above problem.
**Method (definition):** The detailed definition of the method.
**Usage:** How the method is specifically applied to address the problem.
## Scientific Paper to Be Extracted
Title: {title}
Abstract: {abstract}
## Requirements
* Please do not include any method-specific information in problem extraction.
* Similarly, please do not include any problem-specific information in method extraction. The extracted method description (keyword/keyphrase and definition) should be generic and can be applied across all application domains.
* Please do not use abbreviations as keywords/keyphrases.
Please output the extraction results in the JSON format as below. Fields could be "N/A" if no relevant information can be found in the paper title and abstract.
"Problem (keyword/keyphrase)": "…",
"Problem (definition)": "…",
"Problem Discipline": "…",
"Method (keyword/keyphrase)": "…",
"Method (definition)": "…",
"Usage": "…"
Classifying AI4Science
Given the extracted key aspects of papers, we leveraged the following prompt to ask LLM to judge whether the paper solves scientific problems and whether the paper uses AI methods:
## Task Description
Given the extraction results of a research paper, please determine if the main research problem is a scientific problem from traditional disciplines in Science (not including disciplines like Computer Science and Information Science), and if the main method involves the use of Artificial Intelligence.
## Research Paper to Be Extracted
Title: {title}
Abstract: {abstract}
## Extraction Results
results
## Note
* Don’t consider research problem disciplines that involves Computer Science, Information/Data Science as traditional scientific problems.
## Response Format:
Only a dictionary containing the following
"Scientific problem": True/False,
"AI method": True/False,
Model.
For the LLM extraction, we utilized gpt-4o-mini-2024-07-18, which represents a good balance between extraction capability and affordability. We use default generation hyperparameters (e.g., temperature) for all extraction tasks.
Human evaluation.
To evaluate LLM’s ability as an annotator for scholarly papers, we conducted a human evaluation of 50 papers from scientific journals papers and 50 papers from AI conferences, which were randomly selected within our dataset.
Two trained human experts who are familiar with reading scientific literature are assigned to annotate key aspects extraction of each paper, based on the title and abstract, following a carefully designed codebook. They are asked to judge whether the LLM extraction of key aspects is accurate. After that, they are asked whether the paper of interest is an AI4Science paper.
The results are shown in Table 8 9, in which the human agreement is calculated by treating one annotation as the reference and the other as the prediction for each paper. The average extraction accuracy reported by human annotators is 0.910. Together with high human agreement F1 scores (0.876 and 0.946 respectively), this shows that LLM extractions of key aspects are aligned with human preferences. The evaluations for AI4Science classification are also reported, where the F1 score reaches 0.836 for publications from scientific journals. For publications from AI conferences, the F1 score is relatively low (0.647), because of the difficulty in distinguishing research problems in Computer Science domains and general Science domains. However, the high Recall score (0.866) provides a guarantee of avoiding missing important AI4Science work. The difficulty of AI4Science classification is further illustrated by relatively low human agreement(F1 score of 0.632 for both genres), which shows the lack of universally accepted standard of AI4Science. Moreover, this shows that our LLM extraction results are more consistent compared with a group of human annotators.
Extraction acc. Human agreement F1
Scientific pubs. 0.890 0.876
AI pubs. 0.930 0.946
Table 8. Human evaluation of gpt-4o-mini-2024-07-18 extractions.
Classification acc. Human Agreement F1
Prec. Recall F1
Scientific pubs. 0.833 0.840 0.836 0.632
AI pubs. 0.516 0.866 0.647 0.632
Table 9. Human evaluation of gpt-4o-mini-2024-07-18 AI4Science Classification.
Extraction examples.
Listed below are some examples of key aspects extracted by LLM within the human annotation set and whether the paper involves the use of AI methods or tries to solve scientific problems.
Scientific journal paper examples: Please refer to table 10
Title: A machine learning approach to integrate big data for precision medicine in acute myeloid leukemia
Abstract: Cancers that appear pathologically similar often respond differently to the same drug regimens. Methods to better match patients to drugs are in high demand. We demonstrate a promising approach to identify robust molecular markers for targeted treatment of acute myeloid leukemia (AML) by introducing: data from 30 AML patients including genome-wide gene expression profiles and in vitro sensitivity to 160 chemotherapy drugs, a computational method to identify reliable gene expression markers for drug sensitivity by incorporating multi-omic prior information relevant to each gene’s potential to drive cancer. We show that our method outperforms several state-of-the-art approaches in identifying molecular markers replicated in validation data and predicting drug sensitivity accurately. Finally, we identify SMARCA4 as a marker and driver of sensitivity to topoisomerase II inhibitors, mitoxantrone, and etoposide, in AML by showing that cell lines transduced to have high SMARCA4 expression reveal dramatically increased sensitivity to these agents.
LLM Summarized Core Aspects:
Problem (keyword/keyphrase): Patient-Drug Matching in Acute Myeloid Leukemia
Problem (definition): The need for identifying effective drug regimens for patients with acute myeloid leukemia, as similar cancers can respond differently to the same treatments, necessitating a method to better match patients with suitable therapies based on molecular markers.
Problem discipline: Medical Science/Oncology
Method (keyword/keyphrase): Computational Method for Gene Expression Analysis
Method (definition): A computational approach that utilizes data analysis techniques to identify reliable markers from gene expression profiles, potentially incorporating additional omic data to enhance the predictive accuracy for drug sensitivity.
Usage: The method is applied by analyzing gene expression data from acute myeloid leukemia patients to identify molecular markers that are associated with sensitivity to various chemotherapy drugs, thereby improving the precision of treatment matching for individual patients.
AI4Science Classification:
Scientific problem: True
AI method: True
Title: Dynamic microscale flow patterning using electrical modulation of zeta potential
Abstract: Cancers that appear pathologically similar often respond differently to the same drug regimens. Methods to better match patients to drugs are in high demand. We demonstrate a promising approach to identify robust molecular markers for targeted treatment of acute myeloid leukemia (AML) by introducing: data from 30 AML patients including genome-wide gene expression profiles and in vitro sensitivity to 160 chemotherapy drugs, a computational method to identify reliable gene expression markers for drug sensitivity by incorporating multi-omic prior information relevant to each gene’s potential to drive cancer. We show that our method outperforms several state-of-the-art approaches in identifying molecular markers replicated in validation data and predicting drug sensitivity accurately. Finally, we identify SMARCA4 as a marker and driver of sensitivity to topoisomerase II inhibitors, mitoxantrone, and etoposide, in AML by showing that cell lines transduced to have high SMARCA4 expression reveal dramatically increased sensitivity to these agents.
LLM Summarized Core Aspects:
Problem (keyword/keyphrase): Microscale flow control
Problem (definition): The limitations associated with moving fluids at the microscale, particularly the constraints imposed by discrete channels, mechanical valves, and fixed geometries, which restrict the ability to manipulate fluid dynamics effectively.
Problem discipline: Fluid Mechanics
Method (keyword/keyphrase): Field-effect electroosmosis
Method (definition): A technique that utilizes electric fields to induce fluid flow in a medium, allowing for the manipulation of fluid dynamics through the controlled application of electrical signals.
Usage: The method is applied by controlling gate electrodes with an alternating current voltage to generate dynamic flow patterns in a fluidic chamber, enabling real-time modulation of flow characteristics without physical barriers.
AI4Science Classification:
Scientific problem: True
AI method: False
Table 10. Examples of scientific paper extraction and AI4Science judging results
AI conference paper examples: Please refer to table 11
Title: A region-based model for estimating urban air pollution
Abstract: Air pollution has a direct impact to human health, and data-driven air quality models are useful for evaluating population exposure to air pollutants. In this paper, we propose a novel region-based Gaussian process model for estimating urban air pollution dispersion, and applied it to a large dataset of ultrafine particle (UFP) measurements collected from a network of sensors located on several trams in the city of Zurich. We show that compared to existing grid-based models, the region-based model produces better predictions across aggregates of all time scales. The new model is appropriate for many useful user applications such as exposure assessment and anomaly detection.
LLM Summarized Core Aspects:
Problem (keyword/keyphrase): urban air pollution Problem (definition): The challenge of estimating and predicting the dispersion of air pollutants in urban environments, which directly impacts human health and exposure assessments.
Problem discipline: Environmental Science
Method (keyword/keyphrase): Gaussian process model
Method (definition): A statistical method used for regression and classification tasks that relies on the principles of Bayesian inference and provides a flexible way to model complex relationships in data.
Usage: The method is applied to a large dataset of ultrafine particle measurements to estimate urban air pollution dispersion and improve prediction accuracy compared to traditional grid-based models.
AI4Science Classification:
Scientific problem: True
AI method: True
Title: Dependent Relational Gamma Process Models for Longitudinal Networks
Abstract: A probabilistic framework based on the covariate-dependent relational gamma process is developed to analyze relational data arising from longitudinal networks. The proposed framework characterizes networked nodes by nonnegative node-group memberships, which allow each node to belong to multiple latent groups simultaneously, and encodes edge probabilities between each pair of nodes using a Bernoulli Poisson link to the embedded latent space. Within the latent space, our framework models the birth and death dynamics of individual groups via a thinning function. Our framework also captures the evolution of individual node-group memberships over time using gamma Markov processes. Exploiting the recent advances in data augmentation and marginalization techniques, a simple and efficient Gibbs sampler is proposed for posterior computation. Experimental results on a simulation study and three real-world temporal network data sets demonstrate the model’s capability, competitive performance and scalability compared to state-of-the-art methods.
LLM Summarized Core Aspects:
Problem (keyword/keyphrase): Longitudinal networks analysis
Problem (definition): The challenge of analyzing relational data that evolves over time within networks, specifically focusing on the dynamics of node-group memberships and the interactions between nodes in a temporal context
Problem discipline: Statistics and Network Science
Method (keyword/keyphrase): Bayesian inference using Gibbs sampling
Method (definition): A statistical method that involves making inferences about unknown parameters in a model through the use of a sampling technique that generates samples from the posterior distribution of those parameters.
Usage: The method is applied to perform posterior computation for the parameters of the covariate-dependent relational gamma process model, enabling the analysis of the evolving structures of longitudinal networks.
AI4Science Classification:
Scientific problem: False
AI method: True
Table 11. Examples of AI paper LLM extraction and AI4Science judging results
A.3. Semantic Clustering
Embedding
After obtaining the six key aspects of the papers, we first leverage the InstructionEmbedding model(instructor large (Su et al., 2022)) to obtain the semantic embeddings of the scientific problems and AI methods extracted.
For scientific problems, we embed the problem keywords/keyphrases concatenated with the problem definition with the following instructions:
Represent the keyphrase and definition of a scientific problem for clustering and visualizing scientific problems
For AI methods, we embed the AI method keywords/keyphrases concatenated with the AI method definition with the following instructions:
Represent the Artificial Intelligence method paragraph for clustering and visualizing Artificial Intelligence methods
Clustering
After we obtained the semantic embeddings, we used LargeVis (Tang et al., 2016) for dimension reduction followed by HDBSCAN (Campello et al., 2013) for clustering.
Summarization
To have more consistent and generalizable knowledge of the clusters, we asked GPT (gpt-4o-2024-08-06) to summarize each cluster. We first perform TF-IDF on each of the clusters, obtaining a list of top words. Then we provide the top words as well as examples of paper key aspects from a certain cluster to the LLM for summarization.
For scientific problem clusters we use the following prompt:
## Background and Task Description
You are an assistant in scientific research. Given a cluster of scientific problems/challenges, please help summarizing the cluster into a keyword or keyphrase less than 3 words. The top appearing words in the cluster together with some examples of extracted paper from that cluster will be provided. Please summarize based on the provided information.
* Problem (keyword/keyphrase): A keyword or a keyphrase that summarizes the main problem to be addressed in this paper.
* Problem (definition): The detailed definition of the problem.
* Problem discipline: The discipline in which the main problem best fits.
## Top words from texual information
Below are some top words in we extracted from texual information of this cluster using TF-IDF
{top words}
## Examples of scientific papers
Below are some specific examples of scientific challenges in this cluster. For each sample the problem keyword, definition and discipline will be provided.
{examples}
## Summarization Requirement
Based on the above information, please find a keyword or a keyphrase less than *three* words to summarize this cluster of scientific problems, satisfying the following requirements:
* The summarization should be able to cover all the given top words and examples (i.e., with a high "recall")
* The summarization should be specific enough so that it only covers contents in this cluster (i.e., with a high "precision")
* If the summarization is not possible due to lack of information, output "N/A" when applicable.
## Response Format
Please output the summarization as a list, which has exactly *one* elements. The output should be in the format of a python list.
["Keyword"]
For AI method clusters we use the following prompt:
## Background and Task Description
You are an assistant in Artificial Intelligence research, where where AI generally refers to intelligence exhibited by machines (particularly computer systems), including models, algorithms. Given a cluster of Artificial Intelligence methods, please help summarizing the cluster into a keyword or keyphrase less than 3 words. The top appearing words in the cluster together with some examples of Artificial Intelligence metho from that cluster will be provided. Please summarize based on the provided information.
* Method (keyword/keyphrase): A keyword or a keyphrase that summarizes the main method used in this paper to address the above problem.
* Method (definition): The detailed definition of the method.
## Top words from texual information
Below are some top words in we extracted from texual information of this cluster using TF-IDF
{top words}
## Examples of Artificial Intelligence methods
Below are some specific examples of Artificial Intelligence methods in this cluster. For each sample the problem keyword, definition and discipline will be provided.
{examples}
## Summarization Requirement
Based on the above information, please find a keyword or a keyphrase less than *three* words to summarize this cluster of AI methods, satisfying the following requirements:
* The summarization should be able to cover all the given top words and examples (i.e., with a high "recall")
* The summarization should be specific enough so that it only covers contents in this cluster (i.e., with a high "precision")
* If the summarization is not possible due to lack of information, output "N/A" when applicable.
## Response Format
Please output the summarization as a list, which has exactly *one* elements. The output should be in the format of a python list.
["Keyword"]
Merging Clusters
To refine previous clustering results and obtain more consistent clusters, We perform cluster merging based on LLM summarization. If LLM gives the same summary for a group of adjacent clusters on the 2D projection map, we merge them into a large cluster with the same summary. After merging, we have $369$ scientific problems clusters and $345$ AI methods clusters as the final clusters for further analysis.
Appendix B Discrepancy Analysis
Bipartite graph node degrees.
Table 12 presents the nodes with the highest degrees in the bipartite graph, representing the most connected scientific challenges and AI methods in the AI4Science landscape. These high-degree nodes act as key “hubs,” indicating their central role in bridging AI techniques with scientific problems
Discrepancies between the AI and science communities.
Fig. 4-5 illustrates the distribution of AI4Science and non-AI4Science publications across different communities, highlighting how AI and scientific fields engage with interdisciplinary research distinctively. Table.
Figure 4. Distribution of scientific problems of all publications, annotated with gpt-4o-2024-08-06 summary.
Figure 5. Distribution of AI methods of all publications, annotated with gpt-4o-2024-08-06 summary.
(a) Distribution of scientific problems of scientific journal papers, annotated with gpt-4o-2024-08-06 summary.
(b) Distribution of scientific problems of AI conference papers, annotated with gpt-4o-2024-08-06 summary.
(c) Distribution of AI methods of scientific journal papers, annotated with gpt-4o-2024-08-06 summary.
(d) Distribution of AI methods of AI conference papers, annotated with gpt-4o-2024-08-06 summary.
Figure 6. Distribution of scientific problems and AI methods, annotated with gpt-4o-2024-08-06 summary.
Table 13 lists the top clusters emphasized by each community, revealing distinct focus areas and approaches in integrating AI into scientific research.
Scientific Problem Node Deg. AI Method Node Deg.
1 Neural Network Challenges 229 Machine Learning 218
2 Statistical Inference 162 Deep Learning 149
3 Language and Auditory Processing 149 Genomic Analysis 122
4 Optimization Problems 135 Advanced Imaging and Spectroscopy 91
5 Social Media Dynamics 134 Computational Biology 79
6 Decision-Making Processes 126 Protein Design 76
7 Market Economics 118 Neural Networks 70
8 Game Theory 115 Molecular Modeling 65
9 Dynamical Systems 115 Image Processing 64
10 Imaging Challenges 106 Classification Methods 60
11 Causal Inference 105 Data Analysis 58
12 Visual and Sensory Perception 97 Bayesian Inference 57
13 Robotic Actuators 92 Materials Informatics 55
14 Urban and Sustainability Challenges 84 Unsupervised Learning 52
15 Optical and Wireless Communication 79 Wearable Sensor Technology 51
16 Complex Network Dynamics 78 Predictive Modeling 51
17 Scientific Policy Challenges 76 Drug Screening 49
18 Neural Imaging 72 Deep Neural Networks 49
19 Traffic Prediction 69 Convolutional Neural Networks 46
20 Protein Structure and Design 63 Artificial Intelligence Techniques 44
Table 12. Nodes in the bipartite graph (Fig. 2a) with the highest degrees.
Top scientific problems in AI4Science research Top AI methods in AI4Science research
In the science community In the AI community In the science community In the AI community
1 Neural Network Challenges Neural Network Challenges Machine Learning Deep Learning
2 Imaging Challenges Social Media Dynamics Deep Learning Machine Learning
3 Language and Auditory Processing Statistical Inference Genomic Analysis Data Analysis
4 Visual and Sensory Perception Language and Auditory Processing Advanced Imaging and Spectroscopy Neural Networks
5 Protein Structure and Design Game Theory Protein Design Game Theory
6 Materials Design Market Economics Neural Networks Bayesian Inference
7 Artificial Intelligence Ethics Optimization Problems Computational Biology Causal Inference
8 Statistical Inference Decision-Making Processes Molecular Modeling Probabilistic Modeling
9 Single-Cell RNA Sequencing Dynamical Systems Optical Computing Reinforcement Learning
10 Single-Cell Omics Causal Inference Materials Informatics Optimization Methods
11 Decision-Making Processes Robotic Actuators Image Processing Predictive Modeling
12 Cancer Detection Imaging Challenges Neuroimaging Techniques Sequential Decision Making
13 Gene Expression Regulation Optical and Wireless Communication Memory-augmented Systems Convolutional Neural Networks
14 Genetic Variants Traffic Prediction Drug Screening Spatio-temporal Analysis
15 Dynamical Systems Urban and Sustainability Challenges Wearable Sensor Technology Fair Resource Allocation
16 Chemical Reaction Dynamics Fair Resource Allocation Unsupervised Learning Regret Minimization
17 Protein-Protein Interactions Voting and Elections Molecular Dynamics Policy Optimization
18 Wearable Tactile Sensors Visual and Sensory Perception Classification Methods Auction Mechanism Design
19 Antibiotic Resistance Scientific Policy Challenges Deep Neural Networks Classification Methods
20 Motor Control and Learning Complex Network Dynamics Bayesian Inference Generative Models
Table 13. Top scientific problem and AI method clusters in AI4Science research, identified within both the science and AI communities. The clusters are ranked based on the number of AI4Science publications contributed by each community.
Appendix C Experiment Details
C.1. Experiment Setups
Data.
Table 14 lists the statistics of the training and test data splits.
Train Test
Publication year range 2014-2022 2023-2024
Number of All publications 138,577 20,718
AI4Science publications 5,829 1,195
Pub. in well-investigated sci. clusters 3,503 512
Pub. in under-investigated sci. clusters 1,129 379
Pub. in well-investigated AI clusters 3,038 687
Pub. in under-investigated AI clusters 1,336 186
Avg. deg. of Scientific problem nodes 14.4 3.7
Well-investigated sci. nodes 59.5 8.9
Under-investigated sci. nodes 5.3 2.5
AI method nodes 14.7 4.7
Well-investigated AI nodes 25.6 7.4
Under-investigated AI nodes 7.8 1.8
Table 14. The training and test splits of data.
Models.
We examine two approaches to link prediction based on different data representations:
  • Bipartite graph link prediction (“$P_{i}\to M_{i}$” or “$M_{i}\to P_{i}$”): In this approach, scientific problems and AI methods are represented as high-level cluster labels, serving as nodes in a bipartite graph (as defined in Sec. LABEL:). This captures more generalized connections between similar scientific problems and AI methods.
  • Generative link prediction (“$p_{i}\to m_{i},u_{i}$” or “$m_{i}\to p_{i},u_{i}$”): Here, scientific problems and AI methods are expressed in detailed text descriptions rather than coarse cluster labels. This approach leverages textual generative models, such as large language models (LLMs), to capture more fine-grained distinctions and predict links based on textual data.
Following Özer et al. (2024), we employ three categories of bipartite graph link prediction models: (1) link score-based predictions, where we utilize the Katz index calculated using common neighbors (Katz, 1953); and (2) embedding-based link predictions (Perozzi et al., 2014; Tang et al., 2015), in which we apply spectral embedding and node2vec embedding techniques (Grover and Leskovec, 2016).
The Katz index between two nodes $x$ and $y$ are calculated as below:
(1) $\text{Katz}(x,y)=\sum_{\emph{l=1}}^{\infty}\alpha^{l}(A^{l})_{xy},$
where $A$ is the connectivity matrix and $\alpha$ is a penalty factor for path length (In all of our experiments we chose $\alpha=0.1$).
For generative link prediction, we employ two OpenAI GPT models, gpt-3.5-turbo-0125 and gpt-4o-2024-08-06, which represent different levels of generative capabilities. It’s important to note that the training data for gpt-3.5-turbo-0125 extends only up to September 2021, meaning there is no risk of data leakage based on our training/test data splitting strategy. However, the training data for gpt-4o-2024-08-06 includes information up to October 2023, which may introduce potential data leakage and could lead to artificially higher performance. We will discuss the impact of this potential data leakage in the results section.
Evaluation metrics.
For bipartite link prediction, we evaluate performance using Recall, Precision, and F1 scores @K, where K represents the number of generated candidate links, as is standard in the literature (Al Hasan et al., 2006; Yang et al., 2015). Particularly, for a source node $s$, supposing in the test data, it is connected to a set of target nodes $T(s)$, and the model generates $K$ predictions $P(s)$. The metrics are calculated as below:
(2) Precision :=|$:=\frac{|\{p\in T(s)\mid p\in P(s)\}|}{K},$
(3) Recall :=|$:=\frac{|\{t\in P(s)\mid t\in T(s)\}|}{|T(s)|}.$
For link generation, we adopt three widely used evaluation metrics for text generation: ROUGE (Lin, 2004), BLEURT (Sellam et al., 2020), and the cosine similarity of text embeddings. The best scores @K are reported for each of these metrics.
For ROUGE, we report rouge-1-f, which is defined as follows (given a reference’s unigrams r and extraction’s unigrams s):
$p$ $=\frac{|s\cap r|}{|s|}$
$r$ $=\frac{|s\cap r|}{|r|}$
$f$ $=\frac{2*p*r}{p+r}$
For BLEURT (Sellam et al., 2020), we use the pre-trained checkpoint BLEURT-20
For instruction embedding, we first calculate reference’s and extraction’s embedding ($r$ and $e$ respectively) via pre-trained checkpoint Instructor-large (Su et al., 2022) and calculate the cosine similarity:
$cos(r,e)=\frac{<r,e>}{||r||\ ||e||}$
LLM generation setup
For generative link prediction, we consider both Sci $\to$ AI and AI $\to$ Sci generations:
  • Sci $\to$ AI. Given the scientific problem description, generate the AI method that can solve the scientific problem.
  • AI $\to$ Sci. Given the AI method description, generate the scientific problem that the AI methods can solve.
To evaluate LLM link generation quality, we consider the following experiments:
  • Imitation @ K: The nearest K descriptions (scientific problems or AI methods) are imitated to be the predictions. Distance is measured by semantic embedding cosine similarity (Appendix A.3).
  • Direct @ K: In this scenario, we directly ask the LLM to generate K predictions given the extracted scientific problem/AI methods description.
  • RAG @ K: We first find $n$ most similar examples (similarity measure by cosine similarity of semantic embedding) and provide these examples to the LLM to generate K predictions given the extracted scientific problem/AI methods descriptions. We experiment with $n=1,3,5$. In the main paper, we report the results with $n=5$.
Prompts
For generative link prediction with LLMs (LLM+RAG), we use the following prompt [^1] For brevity only Sci $\to$ AI RAG prompt is provided:
## Background and Task Description
You are an expert in both science and artificial intelligence (AI), where AI generally refers to intelligence exhibited by machines (particularly computer systems), including models, algorithms, etc. Given a scientific problem, your task is to recommend potential Artificial Intelligence methods that can be used to address this scientific problem. A few examples of relevant papers with similar scientific problems will be provided.
* Problem (keyword/keyphrase): A keyword or a keyphrase that summarizes the main problem to be addressed in this paper.
* Problem (definition): The detailed definition of the problem.
* Problem discipline: The discipline in which the main problem best fits.
* Method (keyword/keyphrase): A keyword or a keyphrase that summarizes the main method used in this paper to address the above problem.
* Method (definition): The detailed definition of the method.
* Usage: How the method is specifically applied to address the problem.
## Scientific Problem
Please recommend an AI method to address the below scientific problem for writing an academic paper:
{Key Aspects Extraction}
## Examples of AI usage in similar scientific papers: {examples}
## Notes
* If no potential AI method can be used, mark it as "N/A" (not applicable).
* Please respond with specific AI methods instead of high-level AI methods.
* Exactly output one recommendation.
## Response Format
Please output the recommendation of AI methods as a list, which has exactly *one* element. The output should be in the format of a list as below:
"AI Method (keyword/keyphrase)": "…",
"AI Usage": "…"
For graph-level link prediction with LLMs (LLM+Graph), we use the following prompt:  [^2] For brevity only Sci $\to$ AI RAG prompt is provided:
## Background and Task Description
You are an expert in both science and artificial intelligence (AI), where AI generally refers to intelligence exhibited by machines (particularly computer systems), including models, algorithms, etc. Given a scientific problem domain and past usage of Artificial Intelligence methods in solving scientific problems, your task is to recommend potential Artificial Intelligence methods that can be used to address the scientific problem.
## Scientific problem domain
{sci cluster}
## Possible Artificial Intelligence domains
{AI clusters}
## Format of past usage of AI methods to solve Scientific problems:
(u,v,k): Scientific problem u has been solved with AI method v for k times in previous scientific literature.
## Past usage of AI methods to solve Scientific problems:
{example links}
## Notes
* If no potential AI method can be used, mark it as "N/A" (not applicable).
* The AI methods recommended should be within the possible Artificial Intelligence domains.
* The AI methods recommended may or may not be within the given observed links.
* Exactly recommend {k} AI methods.
## Response Format
please output the recommended AI methods as a list, which contains exactly {k} elements. The output should be in the list format as below:
"Artificial Intelligence Method 1",
C.2. Experiment Results
Tables 15-19 summarize the detailed LLM link generation experiment results with different models and evaluation metrics. These results show that:
  • Including RAG examples helps the model make more informed predictions, as all evaluation metric results improve when we increase the examples given from $n=1$ to $n=5$.
  • While Imitation can achieve higher rouge scores, it has less semantic similarity compared with LLM generations.
Sci $\to$ AI ROUGE-1-F BLEURT CosineSim
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Imitation 0.254 0.297 0.313 0.330 0.365 0.397 0.407 0.420 0.853 0.871 0.877 0.883
Direct 0.263 0.286 0.288 0.292 0.407 0.428 0.429 0.429 0.889 0.895 0.895 0.896
RAG n = 1 0.276 0.291 0.300 0.309 0.416 0.434 0.439 0.444 0.888 0.895 0.898 0.901
RAG n = 3 0.286 0.296 0.304 0.313 0.417 0.434 0.439 0.443 0.888 0.895 0.897 0.901
RAG n = 5 0.290 0.297 0.305 0.314 0.417 0.432 0.440 0.444 0.887 0.894 0.897 0.900
Table 15. gpt-3.5-turbo-0125 AI method link generation results
Sci $\to$ AI ROUGE-1-F BLEURT CosineSim
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Imitation 0.254 0.297 0.313 0.330 0.365 0.397 0.407 0.420 0.853 0.871 0.877 0.883
Direct 0.245 0.273 0.284 0.286 0.380 0.410 0.429 0.435 0.885 0.891 0.894 0.895
RAG n = 1 0.266 0.283 0.289 0.294 0.391 0.424 0.438 0.443 0.889 0.894 0.897 0.898
RAG n = 3 0.273 0.292 0.295 0.300 0.395 0.426 0.441 0.447 0.890 0.895 0.898 0.899
RAG n = 5 0.277 0.296 0.300 0.304 0.397 0.430 0.442 0.449 0.890 0.896 0.898 0.900
Table 16. gpt-4o-2024-08-06 AI method link generation results
AI $\to$ Sci ROUGE-1-F BLEURT CosineSim
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Imitation 0.241 0.281 0.293 0.310 0.342 0.372 0.379 0.389 0.795 0.819 0.826 0.836
Direct 0.221 0.236 0.242 0.255 0.383 0.394 0.397 0.403 0.822 0.834 0.837 0.846
RAG n = 1 0.248 0.249 0.253 0.265 0.389 0.390 0.398 0.402 0.823 0.832 0.837 0.845
RAG n = 3 0.246 0.257 0.263 0.261 0.388 0.399 0.402 0.404 0.818 0.832 0.839 0.842
RAG n = 5 0.255 0.267 0.266 0.276 0.387 0.395 0.402 0.407 0.818 0.829 0.838 0.845
Table 17. gpt-3.5-turbo-0125 scientific problem link generation results
AI $\to$ Sci ROUGE-1-F BLEURT CosineSim
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Imitation 0.241 0.281 0.293 0.310 0.342 0.372 0.379 0.389 0.795 0.819 0.826 0.836
Direct 0.227 0.251 0.256 0.254 0.352 0.379 0.388 0.392 0.822 0.840 0.845 0.850
RAG n = 1 0.248 0.267 0.274 0.287 0.356 0.386 0.392 0.406 0.820 0.839 0.845 0.853
RAG n = 3 0.243 0.272 0.283 0.288 0.355 0.385 0.397 0.403 0.820 0.836 0.841 0.851
RAG n = 5 0.249 0.279 0.280 0.285 0.357 0.390 0.397 0.400 0.821 0.838 0.844 0.845
Table 18. gpt-4o-2024-08-06 scientific problem link generation results
Link prediction setting ROUGE-1-F BLEURT CosineSim
@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
Sci $\to$ AI All test data 0.290 0.297 0.305 0.314 0.417 0.432 0.440 0.444 0.887 0.894 0.897 0.900
Well-investigated 0.284 0.300 0.308 0.317 0.418 0.439 0.446 0.447 0.882 0.890 0.893 0.885
Under-investigated 0.301 0.316 0.324 0.333 0.429 0.451 0.459 0.460 0.896 0.904 0.906 0.902
AI $\to$ Sci All test data 0.255 0.267 0.266 0.276 0.387 0.395 0.402 0.407 0.818 0.829 0.838 0.845
Well-investigated 0.243 0.251 0.258 0.262 0.383 0.389 0.395 0.399 0.827 0.840 0.848 0.852
Under-investigated 0.246 0.255 0.265 0.264 0.390 0.398 0.406 0.410 0.822 0.833 0.844 0.848
Table 19. LLM+RAG performances on text generation metrics.

### v2

@@
 However, a significant gap still exists between AI and scientific communities, limiting the full potential of AI methods in driving broad scientific discovery.
-Existing efforts in bridging this gap have often relied on qualitative examination of small samples of literature, offering a limited perspective on the broader AI4Science landscape.
+Existing efforts in identifying and bridging this gap have often relied on qualitative examination of small samples of literature, offering a limited perspective on the broader AI4Science landscape.
 In this work, we present a large-scale analysis of the AI4Science literature, starting by using large language models to identify scientific problems and AI methods in publications from top science and AI venues.
-Leveraging this new dataset, we quantitatively highlight key disparities between AI methods and scientific problems in this integrated space, revealing substantial opportunities for deeper AI integration across scientific disciplines.
+Leveraging this new dataset, we quantitatively highlight key disparities between AI methods and scientific problems, revealing substantial opportunities for deeper AI integration across scientific disciplines.
 Furthermore, we explore the potential and challenges of facilitating collaboration between AI and scientific communities through the lens of link prediction.
 Our findings and tools aim to promote more impactful interdisciplinary collaborations and accelerate scientific discovery through deeper and broader AI integration.
+Our code and dataset are available at: https://github.com/charles-pyj/Bridging-AI-and-Science.
 AI for Science, Large-Scale Literature Analysis, Link Prediction
+copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: Applied computingccs: Information systems Web miningccs: Information systems Collaborative and social computing systems and tools
 1. Introduction
@@
 In particular, the developers of AlphaFold (Jumper et al., 2021) were recognized for their groundbreaking work, which is based on the Transformer network and has revolutionized the prediction of protein structures, providing a transformative approach to solving complex biological problems.
-This is just one example of how AI has emerged as a powerful tool to facilitate scientific research across various disciplines (Wang et al., 2023; Leontidis, 2024; Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
-However, despite the transformative potential of using AI for advancing scientific research (AI4Science), a considerable gap persists between AI and scientific communities, hindering the full exploitation of AI for scientific discovery.
-On one side, advanced AI methodologies might remain mysterious or underutilized by scientists; on the other side, AI researchers may lack awareness of the specific challenges and potential applications in scientific domains, missing the opportunities for interdisciplinary collaborations.
-Existing efforts to identify and bridge this gap have predominantly involved qualitative reviews of AI’s application in science (Wang et al., 2023; Leontidis, 2024), particularly in specialized areas like drug discovery and materials science (Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
-Based on qualitative examination of small samples of literature, these reviews often rely on heuristic insights from domain experts to suggest potential uses of AI in solving scientific problems.
+This compellingly exemplifies how AI has emerged as a powerful tool to facilitate scientific research across various disciplines (Wang et al., 2023; Leontidis, 2024; Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
+However, despite the transformative potential of using AI methods for solving scientific problems (AI4Science), a considerable gap persists between AI and scientific communities, hindering the full exploitation of AI for scientific discovery: On the one side, advanced AI methodologies might remain mysterious or underutilized by scientists; On the other side, AI researchers may lack awareness of the specific challenges and potential applications in scientific domains, missing opportunities for interdisciplinary collaborations.
+These challenges invoke a few key research questions: What does the AI4Science landscape look like?
+Are AI methods and scientific problems evenly interconnected, or are some of them underexplored?
+And how could we expand the connections?
+Existing efforts to approach these questions have predominantly involved small-scale, qualitative reviews of AI’s application in science (Wang et al., 2023; Leontidis, 2024), particularly in specialized areas such as drug discovery and material science (Mak et al., 2023; Guo et al., 2021; Xu et al., 2024).
+These qualitative analyses of the literature often rely on heuristic insights from domain experts to suggest potential uses of AI in solving scientific problems.
 While valuable, such focused reviews are limited in providing a comprehensive and diverse perspective on the AI4Science landscape.
-While recent research from the Science of Science community has conducted larger-scale studies that quantify AI’s impact on publications in scientific fields (Gao and Wang, 2023; Duede et al., 2024), they often heavily rely on established scientific taxonomies in “AI-heavy” fields.
+In the mean time, recent research from the Science of Science community has initialized larger-scale studies that quantify AI’s impact on publications in scientific fields (Gao and Wang, 2023; Duede et al., 2024), although they often heavily rely on established scientific taxonomies in “AI-heavy” fields.
 There remains a lack of delivering a holistic, dynamic, and data-driven overview of AI4Science.
 Such analysis is crucial for understanding the barriers and identifying new opportunities for deep engagement of fast-evolving AI methodology in scientific research.
-In this work, we aim to depict the AI4Science landscape via a large-scale and comprehensive analysis of relevant literature in both scientific and AI communities over the past decade.
+Aiming at addressing these key questions, we conduct a large-scale and comprehensive analysis of relevant literature over the past decade.
 We start by using large language models (LLMs) to identify the scientific problems and AI methods addressed in publications from top science and AI venues, through which we assemble a novel and balanced AI4Science dataset to analyze the role of AI in scientific research (Sec.
+2).
+Leveraging this dataset, we illustrate the AI4Science landscape through two projection maps (Fig.
+2) and a bipartite graph (Fig.
+4a).
+We further quantitatively reveal the current discrepancies between AI and science, leading to several key findings and implications (Sec.
 3).
-Then with projection maps (Fig.
-1) and a bipartite graph (Fig.
-2) constructed from this dataset, we quantitatively reveal the discrepancies between AI and science, leading to several key findings and implications (Sec.
+Finally, through the lens of link prediction, we demonstrate the potential and challenges in advancing the bridge between AI and science (Sec.
 4).
-Finally, through the lens of link prediction, we demonstrate the potential and challenges in advancing the bridge between AI and science (Sec.
-5).
-We anticipate that our findings and tools will foster more effective interdisciplinary collaborations, ultimately accelerating scientific discovery through a more diverse and deeper integration of AI methods.
+Our findings and tools may foster more effective interdisciplinary collaborations and accelerate scientific discovery through a more diverse and deeper integration of AI methods.
 Figure 1.
-Illustration of the AI4Science dataset: (1) LLM-based extraction of scientific problem $p_{i}$, AI method $m_{i}$, and AI usage $u_{i}$ from an example publication.
-(2) 2D projections of the extracted scientific problems $\{p_{i}\}$ and AI methods $\{m_{i}\}$.
-Each dot represents a publication, with coordinates obtained by reducing the dimensionality of the problem/method embeddings.
-(a) Green dots correspond to AI4Science work, where AI is utilized to address scientific problems; Orange dots are papers addressing scientific problems with non-AI techniques; (b) Purple dots indicate using AI to address non-scientific problems.
-The visualizations reveal a noticeable discrepancy in the distribution of AI4Science work (green) versus non-AI4Science work (orange and purple) in both the problem and method spaces.
-Figure 2.
-The AI-Science bipartite graph and node distributions.
-(a) The bipartite graph with scientific problem clusters and AI method clusters as nodes.
-The size of each node corresponds to its unweighted degree, representing the number of AI methods applied to a scientific problem, or the number of scientific problems that applied an AI method.
-For clarity, we omit edges representing fewer than five publications, and isolated nodes are removed thereafter.
-The distribution of links indicates an imbalanced connectivity.
-(b-c) The degree of AI method nodes follows a log-normal distribution, and the degree distribution of scientific problem nodes is even more heavily tailed, suggesting the existence of “hubs” in linking AI and science.
-Notation Value
-Publication year range 2014–2024
-# of publications $N$ 159,295
-# of scientific problem extractions |$|\{p_{i}\}|$ 123,773
-# of scientific problem clusters |$|\{P_{i}\}|$ 369
-# of AI method extractions |$|\{m_{i}\}|$ 35,522
-# of AI method clusters |$|\{M_{i}\}|$ 345
-# of AI usage extractions |$|\{u_{i}\}|$ 35,522
-# of AI4Science publications 7,024
-Table 1. AI4Science dataset $\mathcal{D}=\{(p_{i},P_{i},m_{i},M_{i},u_{i})\}_{i=1}^{N}$.
-2. Related Work
-2.1. Qualitative Literature Reviews of AI4Science
-To identify potential applications of AI techniques in science, researchers typically turn to qualitative literature reviews of relevant papers.
-These reviews include general explorations of AI’s usage across various scientific domains (Wang et al., 2023; Leontidis, 2024), as well as more focused examinations in specific fields such as drug discovery (Dara et al., 2022; Mak et al., 2023; Blanco-Gonzalez et al., 2023; Qureshi et al., 2023), materials design (Sha et al., 2020; Li et al., 2020; Guo et al., 2021), and social sciences (Xu et al., 2024).
-While these reviews provide valuable insights, they often rely on interdisciplinary experts to manually identify and categorize scientific problems and potential AI methods, limiting their scope.
-Such an approach often could not offer a comprehensive and data-driven overview of the broader “AI4Science” landscape.
-2.2. Quantitative Literature Analysis
-In addition to heuristic and qualitative literature reviews, a few recent papers from the Science of Science community have focused on quantitatively analyzing the AI4Science publications.
-For instance, Gao and Wang (2023) examined how AI benefits scientific research, finding that AI use is prevalent across various scientific disciplines, and that papers incorporating AI tend to receive a citation advantage.
-Similarly, Duede et al.
-(2024) explored AI’s engagement in different fields, uncovering an “oil-and-water” phenomenon – while AI-engaged work is spreading across disciplines, it does not integrate well with non-AI-engaged work within each specific field.
-Although these studies offer valuable insights into AI4Science, they primarily focus on areas where AI and science are already “blended” effectively, and overlook the broader and underexplored parts of the landscape.
-As a result, these approaches often fail to offer insights into the disparities between AI and scientific communities.
-In contrast, our work offers a more comprehensive view of the AI4Science landscape by showing and quantifying the connections and gaps between scientific problems and AI methods, as well as insights to uncover potential collaborations.
-3. A New Dataset of AI4Science
-A major challenge in understanding the disparities and potential of AI4Science is the lack of a comprehensive, balanced dataset.
-Existing literature data are often biased toward success stories, such as areas where AI and science are already deeply intertwined.
-Or they reflect the perspective of either the scientific or AI communities, but not both.
-For example, the AI methods discussed in the scientific literature may not represent the full range of advancements in AI, while the scientific applications explored in AI literature may fail to capture the nuances and challenges of real-world scientific problems.
-We take the initiative to create a broad and balanced literature dataset to provide a wide-angled lens for understanding AI4Science research.
+Illustration of LLM-based extraction of scientific problem $p_{i}$, AI method $m_{i}$, and AI usage $u_{i}$ from an example publication.
+The extractions are then semantically projected, clustered, and labeled to form the AI4Science landscape (Fig.
+2), as well as to construct the bipartite graph (Fig.
+4a).
+2. A New Dataset and the AI4Science Landscape
+A major challenge in understanding the disparities and potential of AI4Science is the lack of a comprehensive and balanced dataset.
+Existing literature data are often biased toward successful stories, such as areas where AI and science are already deeply intertwined.
+Or they reflect the perspective of either the scientific or the AI communities, but not both.
+We take the initiative to create a broad and balanced large-scale literature dataset that provides a wide-angled lens for understanding AI4Science research.
 2.1. Curation of AI4Science Literature
 We collect publications from leading science and AI venues to offer a balanced view from both communities:
-  • For science domains, we include three top multidisciplinary journals that represent cutting-edge and high-quality research (Nature, Science, and PNAS) and two of their subjournals (Nature Communications, and Science Advances).
-  • For AI communities, we include six top conferences from the list of CSRankings.org, comprising two AI-focused venues (AAAI, IJCAI), three machine learning venues (ICLR, ICML, NeurIPS), and the Web Conference (WWW) as a venue of applied AI research.
-Our data spans the past decade, from 2014 to 2024, including a total of 159,295 publications from both AI and science venues. More details of the data curation process can be found in Appendix A.1.
+  • For science domains, we include three top multidisciplinary journals that represent cutting-edge and high-quality research (Nature, Science, and PNAS) and two of their subjournals (Nature Communications and Science Advances).
+  • For AI communities, we include seven top conferences from the list of CSRankings.org, comprising two AI-focused venues (AAAI, IJCAI), three machine learning venues (ICLR, ICML, NeurIPS), and two venues of applied AI research (SIGKDD, WWW).
+Our data includes a total of 162,656 publications from 2014 to 2024. Appendix A.1 provides details of the data curation process.
+Figure 2.
+2D projection maps of the AI4Science landscape: (a) the extracted scientific problems $\{p_{i}\}$ and (b) AI methods $\{m_{i}\}$.
+Each dot represents a publication, with coordinates obtained by reducing the dimensionality of the problem/method embeddings.
+Green dots correspond to AI4Science work, using AI methods to address scientific problems; Orange dots show papers addressing scientific problems without using AI methods; Purple dots are papers using AI to address non-scientific problems.
+The visualizations reveal a noticeable discrepancy in the distribution of AI4Science work (green) versus non-AI4Science work (orange and purple) in both the problem and method spaces.
 2.2. Extracting Scientific Problems and AI Methods with LLMs
 To understand how AI is applied to science, we need to extract paired entries of scientific problems and AI solutions from relevant publications.
-Traditionally, this is accomplished through literature reviews and surveys which involve manual examinations of small samples of papers (Wang et al., 2023; Leontidis, 2024), a labor-intensive and time-consuming process.
+Traditionally, this is accomplished through literature reviews and surveys which involve manual examinations of small samples of papers (Wang et al., 2023; Leontidis, 2024), a process that is labor-intensive and time-consuming.
 Recent quantitative studies in the Science of Science field have started using metadata entries, such as keywords, to identify AI usage and categorize scientific topics (Gao and Wang, 2023; Duede et al., 2024).
@@
 If a paper addresses a scientific problem using an AI method – i.e., $p_{i}$, $m_{i}$, and $u_{i}$ are all non-empty – it is considered as an AI4Science work.
-Figure 1 presents an example of the extraction results of an AI4Science publication.
+Fig.
+1 presents an example of the extraction results of an AI4Science publication.
 To assess the reliability of GPT extractions, we conducted a small-scale human evaluation on 100 papers.
@@
 For more details on the extraction prompts, additional examples of the extracted data, and verification with human annotations, please refer to Appendix A.2.
-3.3. Semantic Clustering of Scientific Problems and AI methods
-One of the key differences between this new AI4Science dataset and the data used in existing studies (Gao and Wang, 2023; Duede et al., 2024) is that our dataset contains detailed textual descriptions of the identified scientific problems and AI methods rather than pre-defined taxonomies, allowing for a deeper semantic analysis of their relationships.
-To achieve this, we generate semantic embeddings for the scientific problems ${p_{i}}$, AI methods ${m_{i}}$, and AI usages ${u_{i}}$ using the InstructorEmbedding model (Su et al., 2022).
+2.3. Semantic Projection and Clustering
+One of the key advantages of this new AI4Science dataset compared to the data used in existing studies (Gao and Wang, 2023; Duede et al., 2024) is that our dataset contains detailed textual descriptions of the identified scientific problems and AI methods rather than pre-defined taxonomies, allowing for a deeper semantic analysis of their relationships.
+To achieve this, we generate semantic embeddings for the scientific problems $\{p_{i}\}$ and AI methods $\{m_{i}\}$ using the InstructorEmbedding model (Su et al., 2022).
 We then apply LargeVis (Tang et al., 2016) to project these high-dimensional embeddings into 2D coordinates.
-Next, HDBSCAN (Campello et al., 2013) is utilized to group similar scientific problems and AI methods.
-For each cluster, we use GPT-4o to summarize the main topic based on frequent keywords and samples of extraction results.
-This process results in 369 and 345 clusters respectively, with corresponding cluster labels $P_{i}$ and $M_{i}$ to each data point.
-Fig.
-1 displays the 2D visualizations, and Fig.
-4-5 in the Appendix presents the maps with cluster names annotated.
-Details of the process can be found in Appendix A.3.
-Unlike directly using keywords or categorizations from publication metadata (Gao and Wang, 2023; Duede et al., 2024), our approach employs a data-driven methodology.
-By using clustering and summarization, relevant AI methods discussed in both scientific and AI literature are grouped together, as are the scientific problems mentioned in both fields.
-This ensures greater consistency and generalizability of the data entries.
+Next, a density-based clustering method, HDBSCAN (Campello et al., 2013) is utilized to group similar scientific problems and AI methods in this 2D space.
+For each cluster, we use GPT-4o to generate an understandable summarization based on frequent keywords and samples of extraction results.
+This process results in 390 and 355 clusters respectively, with corresponding cluster labels $P_{i}$ and $M_{i}$ to each data point.
+The choice of the methods and the details of the process are in Appendix A.3.
+Fig. 2 shows the 2D maps that illustrate the AI4Science landscape. Fig. 5-6 in Appendix present the maps with the cluster labels.
+Unlike using keywords or categorizations from publication metadata (Gao and Wang, 2023; Duede et al., 2024), our approach employs a data-driven methodology.
+By using textual embedding, clustering, and topic labeling, relevant AI methods discussed in both scientific and AI literature are grouped together, as are the scientific problems mentioned in both fields.
+This enables a semantic understanding of the AI4Science space and ensures greater consistency and generalizability of the data.
 2.4. Dataset Overview and the Bipartite Graph
@@
 To better illustrate the connections between AI and science, we construct a bipartite graph based on the clusters.
-In this bipartite graph, scientific problem clusters and AI method clusters represent two types of nodes, and the publications act as edges connecting them.
+In this bipartite graph, scientific problem clusters and AI method clusters are represented as two types of nodes, and the publications act as edges connecting them.
 Formally, the graph is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}=\{P_{i}\}\cup\{M_{i}\}$ and $\mathcal{E}=\mathcal{D}$.
 As shown in Fig.
-2(a), the graph provides an intuitive visual representation of how AI is linked to scientific challenges and vice versa.
+4a, the graph provides an intuitive visual representation of how AI is linked to scientific challenges and vice versa.
+Notation Value
+Publication year range 2014–2024
+# of venues 12
+# of publications $N$ 162,656
+# of scientific problem extractions |$|\{p_{i}\}|$ 129,038
+# of scientific problem clusters |$|\{P_{i}\}|$ 390
+# of AI method extractions |$|\{m_{i}\}|$ 42,455
+# of AI method clusters |$|\{M_{i}\}|$ 355
+# of AI usage extractions |$|\{u_{i}\}|$ 42,455
+# of AI4Science publications 7,542
+Table 1. AI4Science dataset $\mathcal{D}=\{(p_{i},P_{i},m_{i},M_{i},u_{i})\}_{i=1}^{N}$.
 (a) Scientific problem clusters.
@@
 Figure 3.
-Scientific problem clusters (a) and AI method clusters (b) are visualized as circles in the scatter plots.
+Scientific problem clusters (a) and AI method clusters (b) are visualized as scatters.
 In both plots, the $x$-axis represents the total number of publications in each cluster, while the $y$-axis reflects the number of interdisciplinary AI4Science publications in each cluster.
-Three outstanding points are skipped for clarity1.
 The black lines show the regression results on the clusters.
-Clusters falling below the regression line highlight the under-investigated regions, where the integration of AI and science remains limited.
+Clusters above the line indicate regions well-explored for AI4Science.
+Clusters falling below the line highlight the under-explored regions, where the integration of AI and science remains limited.
 These areas represent potential opportunities for further interdisciplinary collaboration.
 3. Discrepancy Between AI and Science
-Both Fig. 1 and 2 show a noticeable pattern of discrepancy between AI and science. This section quantitatively assesses these disparities, providing a deeper understanding of the gaps.
+Both Fig. 2 and 4a show a noticeable pattern of discrepancy between AI and science. This section quantitatively assesses these disparities, providing a deeper understanding of the gaps.
 3.1. Uneven Distribution of AI4Science Research
@@
 2 illustrate the distribution of AI4Science work within the semantic spaces of scientific problems and AI methods (the green dots among orange and purple dots).
-The uneven distributions suggest that AI4Science work is heavily clustered in certain subareas, while sit remains underrepresented in wide ranges of both the problem and the method spaces.
+The uneven distributions suggest that AI4Science work is heavily clustered in certain subareas, while it remains underrepresented in wide ranges of both the problem and the method spaces.
 In Fig.
@@
 The slopes of the regression lines represent the average proportions of AI4Science work in the literature of each scientific problem cluster or AI method cluster.
-Clusters below the regression line indicate problems or AI methods in which the integration of AI4Science is underrepresented (noted the under-investigated regions), suggesting potential opportunities for further interdisciplinary collaboration.
-Under-investigated scientific problems.
+Clusters below the regression line indicate problems or AI methods in which the integration of AI4Science is underrepresented (noted the under-explored regions), suggesting potential opportunities for further interdisciplinary collaboration.
+11footnotetext: Six outstanding data points are skipped from the scatter plots for visualization clarity: ‘Neural Network Challenges’ (820,375) and ‘Urban Traffic Management’ (551,298) in (a), ‘Machine Learning’ (770,539), ‘Deep Learning Models’ (605,344), ‘Genomic Analysis Methods’ (292,287), and ‘Molecular Dynamics Simulations’ (214,210) in (b).
+Underexplored scientific problems.
 In the bottom-right section of Fig.
-3(a), below the regression line, several scientific problem clusters are highlighted as being rarely addressed by AI techniques.
-Notable examples include Magnetic Materials, Asymmetric Synthesis, Lithium-Ion Battery Challenges, Antibiotic Resistance, Stem Cell Differentiation, Viral Mechanisms, Climate Change Emissions, and Enzyme Catalysis.
-These scientific problems could benefit from more engagement of AI methods.
-Under-investigated AI methods.
+3a, below the regression line, several scientific problem clusters are highlighted as being rarely approached through AI techniques.
+Notable examples include ‘Asymmetric Synthesis’, ‘Magnetic Materials and Spintronics’, ‘Forests and Climate Change’, ‘Viral Mechanisms’, ‘Evolutionary Paleontology’, ‘Nanostructure Synthesis’, ‘Diabetes and Insulin Regulation’, ‘Bacterial Infections and Immune Response’, ‘Cancer Signaling Pathways’, and ‘Cancer Drug Resistance’.
+Scientific problems in this region could benefit from more engagement of AI methods.
+Underutilized AI methods.
 Similarly, Fig.
-3(b) reveals AI techniques that have not yet been widely applied to sciences, including: Attention Mechanism, Gradient-Based Methods/Gradient Descent, Embedding Techniques, Policy Optimization, Regularization Techniques, Theoretical Analysis, Adversarial Training, Contrastive Learning, and Variational Inference.
-These techniques could have been explored in a wider range of scientific applications.
-11footnotetext: Three outstanding data points are skipped from the scatter plots for visualization clarity: Neural Network Challenges (838, 653) in (a), Machine Learning (601, 446) and Deep Learning (577, 308) in (b).
-Node type Partition Avg. Deg.
-Scientific problem clusters (n=369) All nodes 15.7
-Well-investigated 63.2
-Under-investigated 6.3
-AI method clusters (n=345) All nodes 16.5
-Well-investigated 29.2
-Under-investigated 8.5
-Table 2. Average degree of nodes in the bipartite graph.
+3b reveals AI techniques that have not yet been widely applied to sciences, including: ‘Attention Mechanisms’, ‘Gradient-Based Methods’, ‘Graph Embedding’, ‘Regularization Methods’, ‘Transformer Architectures’, and ‘Contrastive Learning’.
+These techniques could have been utilized in a wider range of scientific applications.
+Figure 4.
+The AI-Science bipartite graph and node distributions.
+(a) The bipartite graph with scientific problem clusters and AI method clusters as nodes.
+The size of each node corresponds to its unweighted degree, representing the number of AI methods applied to a scientific problem, or the number of scientific problems utilizing an AI method.
+For visualization clarity, edges representing fewer than four publications and the resulting isolated noes are hidden.
+The distribution of links indicates imbalanced connectivity.
+(b-c) The degree of AI method nodes follows a log-normal distribution, and the degree distribution of scientific problem nodes is even more heavily tailed, suggesting the existence of “hubs” in linking AI and science.
 3.2. Uneven Distribution of Bipartite Links
@@
 Quantitatively, heavy-tailed degree distributions are observed as visualized in Fig.
-2(b-c).
-Specifically, the degrees for AI method nodes follow a log-normal distribution, while the degree distribution for scientific problem nodes exhibits an even heavier tail, though not as extreme as a power-law distribution.
+4(b-c).
+Specifically, the degrees for AI method nodes follow a log-normal-like distribution, while the degree distribution for scientific problem nodes exhibits an even heavier tail, though not as extreme as a power-law distribution.
 These findings support the idea that a small number of “hubs” are linked to a large variety of AI methods or scientific problems, while many other nodes remain peripheral to these interdisciplinary connections.
@@
 This suggests that the structure of the AI4Science bipartite graph reflects a general pattern of interdisciplinary and collaborative work.
-Table 12 in the Appendix lists the “hubs” with the highest node degrees, representing the most interconnected scientific problems and AI methods.
-From the table, we observe that the scientific problems linked to the widest variety of AI methods are primarily computational or data-processing problems, such as: Neural Network Challenges, Statistical Inference, Language and Auditory Processing.
-These are followed by more domain-specific problems like: Social Media Dynamics, Decision-Making Processes, Market Economics.
-On the AI methods side, the techniques applied to the greatest number of scientific challenges include general methods such as: Machine Learning, Deep learning, Neural Networks, Classification Methods, Data Analysis, Bayesian Inference.
-In addition, there are specialized techniques tailored to scientific domains, including: Genomic Analysis, Computational Biology, Protein Design, Molecular Modeling, Material Informatics.
+Table 13 in Appendix lists the “hubs” with the highest node degrees, representing the most interconnected scientific problems and AI methods.
+From the table, we observe that the scientific problems linked to the widest variety of AI methods are primarily computational or data-processing problems, such as: ‘Neural Network Challenges’ and ‘Statistical Inference’.
+These are followed by more domain-specific problems like: ‘Urban Traffic Management’, ‘Electronic Health Data Challenges’, ‘Visual Perception and Neuroscience’, ‘Market Pricing and Allocation’, ‘Single-cell RNA Sequencing’, ‘Soft Robotics’, and ‘Drug Discovery’.
+On the AI methods side, the techniques applied to the greatest number of scientific challenges include general methods such as: ‘Machine Learning’, ‘Deep learning Models’, ‘Neural Networks’, ‘Classification Methods’, ‘Data Analysis’.
+In addition, there are specialized techniques tailored to scientific domains, including: ‘Genomic Analysis Methods’, ‘Computational Drug Screening’, ‘Molecular Dynamics Simulations’, ‘Protein Structure Design’, and ‘Molecular Structural Design’.
+Node type Partition Avg. Deg.
+Scientific problem clusters (n=390) All nodes 13.0
+Well-explored 41.3
+Under-explored 5.0
+AI method clusters (n=355) All nodes 14.3
+Well-explored 19.1
+Under-explored 8.1
+Table 2. Average degree of nodes in the bipartite graph.
 We further break down the node degree distribution based on cluster partitions.
-Specifically, we calculate the average node degree for well-investigated and under-investigated clusters, as identified earlier in Sec.
-4.1.
+Specifically, we calculate the average node degree for well-explored and under-explored clusters, as identified earlier in Sec.
+3.1.
 The statistics are listed in Table 2.
-We observe a higher average degree for well-investigated scientific problems and AI methods.
-This suggests that under-investigated regions not only have fewer AI4Science publications, but also exhibit a lesser variety, or limited connectivity in the bipartite graph.
+We observe a higher average degree for well-explored scientific problems and AI methods.
+This suggests that under-explored regions not only have fewer AI4Science publications, but also exhibit a lesser variety, or limited connectivity in the bipartite graph.
 3.3. Discrepancy Between the AI and Science Communities
 Our dataset includes publications from both scientific journals and AI conferences, allowing us to explore how these two communities approach the integration of AI into science in distinct ways.
-Table 13 lists the top clusters emphasized by each community.
-In terms of facilitating scientific discovery with AI, the science community places greater emphasis on challenges such as: Protein Structure and Design, Materials Design, Artificial Intelligence Ethics, and Single-Cell RNA Sequencing; In contrast, the AI community tends to focus more on areas like: Social Media Dynamics, Game Theory, Market Economics, etc.
+Table 14 in Appendix lists the top clusters emphasized by each community.
+In terms of facilitating scientific discovery with AI, the science community places greater emphasis on challenges such as: ‘Single-cell RNA Sequencing’, ‘Mechanical Properties of Materials’, ‘Protein Design and Function Prediction’, ‘Cancer Detection’, and ‘Genetic Variants and Traits’; In contrast, the AI community tends to focus more on areas like: ‘Urban Traffic Management’, ‘Electronic Health Data Challenges’, ‘Statistical Inference’, ‘Market Pricing and Allocation’, ‘Soft Robotics’, etc.
 When it comes to applying AI methods to scientific problems, the two communities also show different preferences.
-In addition to general approaches like Machine Learning and Deep Learning, the science community frequently employs AI techniques specifically tailored for scientific challenges, such as: Genomic Analysis, Protein Design, and Computational Biology.
-In comparison, the AI community utilizes a broader range of general AI methodologies, covering diverse areas such as: Data Analysis, Game Theory, Causal Inference, Probabilistic Modeling, Reinforcement Learning, Optimization Methods, Generative Models, etc.
+In addition to general approaches like ‘Machine Learning’ and ‘Deep Learning Models’, the science community frequently employs AI techniques specifically tailored for scientific challenges, such as: ‘Genomic Analysis Methods’, ‘Molecular Dynamics Simulations’, and ‘Computational Drug Screening’.
+In comparison, the AI community utilizes a broader range of general AI methodologies, covering diverse techniques such as: ‘Reinforcement Learning’, ‘Attention Mechanisms’, ‘Policy Gradient Methods’, ‘Multitask Learning’, ‘Spatio-Temporal Analysis’, ‘Causal Inference’, ‘Self-Supervised Learning’, and ‘Unsupervised Learning’.
 In addition, by comparing Table 14 with Table 13, we also find that the scientific problems addressed by the AI community, tend to be connected to a wider variety of AI methods.
@@
 This highlights the discrepancy between the two communities, where the science community engages with a limited range of AI methods compared to the AI community, and vice versa.
-The distribution of AI4Science work published in scientific journals and AI conferences in the semantic maps (Fig. 6 in the Appendix) also illustrates the different focuses of these two communities.
+The distribution of AI4Science work published in scientific journals and AI conferences in the semantic maps (Fig. 7 in Appendix) also illustrates the different focuses of these two communities.
 3.4. Findings and Implications
@@
   1. (F1)
-    Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods under-investigated in the collaboration context (Sec.
-4.1);
+    Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods underexplored in the collaboration context (Sec.
+3.1);
   2. (F2)
@@
   1. (I1)
-    More attention should be directed toward exploring the under-investigated areas, incorporating a broader range of scientific challenges and AI techniques into the AI4Science landscape.
+    More attention should be directed toward exploring the under-explored areas, incorporating a broader range of scientific challenges and AI techniques into the AI4Science landscape.
   2. (I2)
-    Efforts should be made to discover new connections between AI and science, uncovering innovative ways to apply AI methods to scientific research.
-5. Bridging AI and Science: Through the Lens of Link Prediction
-Building on the findings and implications above, we delve deeper into the potentials and challenges of advancing the connections between AI and science from the perspective of link prediction.
+    Efforts should be made to discover new connections between AI and science, uncovering innovative ways to apply AI methodologies to scientific research.
+In the following section, we present preliminary explorations of addressing these implications through link prediction.
+4. Bridging AI and Science:
+Through the Lens of Link Prediction
+Building on the findings and implications above, we dive deeper into the potentials and challenges of advancing the connections between AI and science from the perspective of link prediction.
 Setting Method Precision Recall F1
 @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
-Sci $\to$ AI Katz index 0.430 0.335 0.320 0.272 0.093 0.223 0.311 0.450 0.153 0.267 0.315 0.339
-Node2vec 0.988 0.836 0.719 0.530 0.259 0.505 0.625 0.765 0.411 0.630 0.669 0.626
-LLM+RAG 0.326 0.216 0.200 0.140 0.058 0.096 0.129 0.150 0.099 0.132 0.157 0.145
-LLM+Graph 0.541 0.594 0.530 0.437 0.137 0.311 0.419 0.558 0.218 0.408 0.468 0.490
-AI $\to$ Sci Katz index 0.485 0.402 0.362 0.291 0.095 0.205 0.262 0.344 0.160 0.271 0.304 0.316
-Node2vec 0.997 0.904 0.837 0.717 0.197 0.371 0.470 0.622 0.329 0.526 0.602 0.666
-LLM+RAG 0.368 0.293 0.267 0.210 0.036 0.057 0.077 0.098 0.066 0.096 0.120 0.133
-LLM+Graph 0.458 0.316 0.300 0.201 0.068 0.113 0.163 0.152 0.118 0.167 0.211 0.173
-Table 3. Link prediction results of different models.
-Method Link prediction setting Precision Recall F1
-@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
-Node2vec Sci $\to$ AI Well-investigated 1.000 0.980 0.922 0.799 0.115 0.295 0.417 0.606 0.207 0.453 0.574 0.689
-Under-investigated 0.960 0.726 0.535 0.310 0.392 0.734 0.841 0.928 0.557 0.730 0.654 0.465
-AI $\to$ Sci Well-investigated 0.996 0.956 0.916 0.807 0.126 0.293 0.414 0.616 0.224 0.448 0.571 0.699
-Under-investigated 0.978 0.651 0.481 0.284 0.560 0.836 0.918 0.978 0.712 0.732 0.631 0.440
-LLM+RAG Sci $\to$ AI Well-investigated 0.422 0.271 0.247 0.189 0.031 0.054 0.074 0.102 0.057 0.090 0.113 0.132
-Under-investigated 0.240 0.178 0.157 0.111 0.081 0.149 0.182 0.224 0.121 0.162 0.169 0.148
-AI $\to$ Sci Well-investigated 0.344 0.259 0.267 0.206 0.027 0.053 0.076 0.098 0.050 0.088 0.118 0.133
-Under-investigated 0.188 0.115 0.089 0.066 0.090 0.140 0.159 0.167 0.122 0.126 0.114 0.095
+Sci $\to$ AI Katz index 0.115 0.108 0.086 0.076 0.015 0.035 0.052 0.104 0.026 0.053 0.065 0.088
+Node2vec 0.198 0.185 0.168 0.129 0.033 0.098 0.138 0.197 0.056 0.128 0.151 0.156
+LLM (Cluster) 0.352 0.300 0.238 0.159 0.053 0.101 0.134 0.176 0.093 0.151 0.171 0.167
+LLM (Paper) 0.282 0.227 0.200 0.156 0.058 0.106 0.129 0.155 0.096 0.145 0.157 0.155
+AI $\to$ Sci Katz index 0.094 0.120 0.110 0.102 0.010 0.028 0.041 0.079 0.018 0.045 0.060 0.089
+Node2vec 0.096 0.108 0.110 0.109 0.013 0.042 0.058 0.104 0.023 0.060 0.076 0.106
+LLM (Cluster) 0.201 0.168 0.143 0.129 0.016 0.042 0.053 0.092 0.030 0.068 0.077 0.107
+LLM (Paper) 0.377 0.323 0.307 0.296 0.065 0.104 0.127 0.163 0.111 0.158 0.180 0.211
+Table 3. Link prediction results of different models. Bold numbers highlights the highest performances in each scenario.
+Link prediction Precision Recall F1
+Setting Method Region @1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
+Sci $\to$ AI LLM (Paper) Well-explored 0.293 0.259 0.230 0.184 0.025 0.059 0.073 0.105 0.046 0.095 0.111 0.134
+Under-explored 0.250 0.195 0.171 0.129 0.111 0.192 0.224 0.247 0.153 0.193 0.194 0.169
+LLM (Cluster) Well-explored 0.360 0.411 0.353 0.228 0.030 0.082 0.123 0.164 0.055 0.137 0.182 0.191
+Under-explored 0.194 0.125 0.089 0.051 0.068 0.122 0.147 0.182 0.101 0.124 0.111 0.079
+AI $\to$ Sci LLM (Paper) Well-explored 0.403 0.340 0.334 0.332 0.045 0.078 0.096 0.126 0.080 0.127 0.149 0.182
+Under-explored 0.364 0.301 0.267 0.236 0.151 0.222 0.264 0.311 0.214 0.255 0.265 0.269
+LLM (Cluster) Well-explored 0.328 0.251 0.183 0.185 0.020 0.047 0.053 0.102 0.038 0.079 0.082 0.132
+Under-explored 0.076 0.057 0.074 0.102 0.025 0.051 0.051 0.037 0.038 0.054 0.060 0.055
 Table 4. Link prediction results on well- and under-explored regions.
@@
 We split the data into training and test sets based on publication dates.
-The training set includes papers published between 2014 and 2022 (5,829 AI4Science publications), which are used for training models, retrieving data points for generation augmentation, and identifying well/under-investigated areas, as illustrated in Fig.
+The training set includes papers published between 2014 and 2022 (6287 AI4Science publications), which are used for training models, retrieving data points for generation augmentation, and identifying well/under-explored areas, as illustrated in Fig.
 3.
-The testing set consists of publications from 2023 and 2024 (1,195 AI4Science publications), serving as the ground truth for evaluating link prediction models.
-The statistics are provided in Table 14 in the Appendix.
-It is important to note that the analyses prior to this section involve the full AI4Science dataset, spanning the period from 2014 to 2024.
+The testing set consists of publications from 2023 and 2024 (1225 AI4Science publications), serving as the ground truth for evaluating link prediction models.
+The statistics are provided in Table 15 in Appendix.
+Note that the analyses prior to this section involve the full AI4Science dataset, spanning the period from 2014 to 2024.
 Model and evaluation.
 We employ two types of models to predict links between scientific problems and AI methods.
-The first category includes conventional bipartite link prediction models (Özer et al., 2024), such as those based on the Katz index (Katz, 1953) and node2vec embeddings (Grover and Leskovec, 2016).
-Formally, given a source node $P_{i}$ (or $M_{i}$) in the bipartite graph, we predict $K$ potential target nodes for linking.
-We then compare the predicted set with the ground truth target node set and report Precision, Recall, and F1 scores @K (Al Hasan et al., 2006; Yang et al., 2015).
-The bipartite graph is constructed and the models are trained using only the training data, with evaluation performed on the testing data.
-Please note that existing links may also be counted in constructing the ground truth set, because a new AI4Science work may reuse an existing connection between a scientific problem and an AI method.
-In addition to conventional link prediction, we utilize a large language model (LLM) for generative predictions.
-Specifically, given a scientific problem $p_{i}$ from cluster $P_{i}$, we generate potential AI methods $m_{i}$ that could be applied.
-These generated texts are then mapped to cluster labels $M_{i}$ through embedding similarity matching.
-Similarly, given an AI method $m_{i}$, we can generate potential scientific problems $p_{i}$ that it could address.
-We use the OpenAI gpt-3.5-turbo-0125 model as the generator.
-Since this model’s knowledge is current up to September 2021, there is no risk of data leakage.
+The first category directly make predictions on the cluster level with conventional bipartite link prediction models (Özer et al., 2024), such as those based on the Katz index (Katz, 1953) and node2vec embeddings (Grover and Leskovec, 2016) or large language models (LLMs).
+Formally, given the cluster-level bipartite graph constructed on the training data and a source node $P_{i}$ (or $M_{i}$), the model predicts $K$ target nodes that could be potentially linked to the source node.
+We then compare the predicted nodes with the ground-truth target node set in the test data and report Precision, Recall, and F1 scores @K (Al Hasan et al., 2006; Yang et al., 2015).
+Note that existing links may also appear in the ground-truth set, because a new AI4Science work may repeat an existing connection between a scientific problem and an AI method.
+Specifically, when using LLM for link prediction, the bipartite links to the source nodes are provided in prompts, and the model is referred to as ‘LLM (Cluster)’.
+In addition, we also utilize LLMs for paper-level generative predictions.
+Specifically, given a scientific problem $p_{i}$ from cluster $P_{i}$, we prompt the LLM to generate potential AI methods $m_{i}$ that could be applied.
+These generated method descriptions are then mapped to cluster labels $M_{i}$ through embedding similarity.
+Similarly, given an AI method $m_{i}$, we prompt the LLM to generate potential scientific problems $p_{i}$ that it could address.
 The prediction process is further enhanced by retrieving similar papers from the training set using retrieval-augmented generation (RAG) (Lewis et al., 2020).
-While we can use large language models for generative predictions leveraging embedding similarity matching, the large language model is only exposed to a small neighborhood of literature through RAG and has no access to the higher-level linked bipartite graph structure.
-Thus, we include results where the entire graph (training links) is provided to the model.
-Given a scientific problem cluster $P_{i}$, all possible AI method clusters $M$ and past links between the clusters, we directly generate AI method cluster labels $M_{i}$.
-Similarly, given a AI methods cluster $M_{i}$, all possible scientific problem clusters $P$ and past links between the clusters, we directly generate scientific problem cluster labels $P_{i}$.
+The model is referred to as ‘LLM (Paper)’.
+We use OpenAI’s gpt-4o-2024-08-06 as the LLM model and report its performance in the main text.
+Additionally, we also tested the gpt-3.5-turbo-0125 model as the generator, given that its knowledge is limited to September 2021, ensuring no risk of data leakage.
+Our results show that the performance of LLM (Paper) remains consistent, suggesting a low risk of data leakage.
+However, the performance of LLM (Cluster) declines, likely due to GPT-3.5’s limited ability of reasoning and processing long contexts.
 For more detailed descriptions of the models and evaluation metrics, please refer to Appendix C.
 4.2. Overall Link Prediction Results
-Table 3 presents models’ performances on the test data, where conventional link prediction methods show strong results in both Precision and Recall.
-Particularly, while both Katz and node2vec rely on suggesting similar nodes to link, node2vec outperforms Katz, suggesting the underlying mechanism connecting scientific problems and AI methods goes beyond homophily in the local neighborhood and potentially considers the broader structures of the AI4Science network.
-Notably, the LLM+RAG approach lags behind conventional link prediction methods, emphasizing the challenge of generative link prediction at the individual paper level.
-A potential explanation is that, unlike node-level link prediction in a bipartite graph, LLMs struggle to obtain a comprehensive global view of the AI4Science landscape, even with the aid of RAG.
-This implies the benefit of considering the global bipartite graph structures and the opportunity to build knowledge graphs in the AI4Science domain to help LLMs better understand the broader context.
-This is further elaborated by the results for LLM+Graph, where the model is provided with global bipartite graph structure.
-LLM+Graph outperforms LLM+RAG, suggesting that LLM benefits from bipartite linked graph information.
-Moreover, LLM+Graph achieves comparable results with Katz index, which implies LLM’s ability to understand graph-level features and make informed link predictions.
-Nevertheless, LLM+RAG offers a unique advantage: it can generate detailed ideas in texts for AI4Science applications.
-Table 19 in the appendix reports evaluation results on text generation metrics, where LLM+RAG demonstrates notable results.
-5.3. Prediction in the Well- and Under-investigated Regions
+Table 3 presents the models’ performances on the test data, where LLM-based link prediction methods outperform conventional approaches in most settings, highlighting their strong potential for predicting AI4Science research directions.
+[…]
+For conventional link prediction methods, while both Katz and node2vec rely on suggesting similar nodes to link to, node2vec outperforms Katz, suggesting the underlying mechanism connecting scientific problems and AI methods goes beyond homophily in the local neighborhood and potentially considers the broader structures of the AI4Science network.
+This finding further highlights the potential utility of the AI4Science bipartite graph.
+4.3. Well- and Under-Explored Regions
 Fig.
-3 reveals the presence of well- and under-investigated areas in both the scientific problem and AI method spaces.
+3 reveals the presence of well- and under-explored areas in both the scientific problem and AI method spaces.
 Echoing Implication (I1), we comparatively explore the potentials and challenges in facilitating interdisciplinary work for these regions.
-Formally, we refer to scientific problem clusters or AI method clusters above the regression lines as well-investigated areas, while those that fall below the lines are the under-investigated regions.
+Formally, we refer to scientific problem clusters or AI method clusters above the regression lines as well-explored areas, while those that fall below the lines are the under-explored regions.
 Clusters that reside within the 95% confidence intervals of the regression results are excluded from this analysis.
-Table 4 compares the link prediction results for well-investigated and under-investigated clusters.
-In both categories of link prediction methods, we observe a similar trend: Precision scores are generally higher for well-investigated scientific problems and AI methods, even as K increases.
-In contrast, Recall is relatively higher for under-investigated clusters.
-These results are consistent with the node degree statistics in Table 2, where the average degree of under-investigated nodes is much lower, making it easier for the models to predict links, resulting in higher Recall.
-For instance, node2vec can retrieve nearly all ground truth links with K=10.
-Additionally, in the under-investigated partition, the rapid decline in Precision as K increases suggests that link prediction models are capable of discovering novel links that have not been identified by human researchers in the test data.
-This highlights great opportunities for expanding the connections between AI methods and scientific challenges.
-Reference Method # of novel (unseen) links
+Note that for this experiment, we use only the publications in the training split to decide the well- and under-explored regions.
+Table 4 compares the link prediction results for well-explored and under-explored clusters.
+In both categories of link prediction methods, we observe a similar trend: Precision scores are generally higher for well-explored scientific problems and AI methods, even as K increases.
+In contrast, Recall is relatively higher for under-explored clusters.
+These results are consistent with the node degree statistics in Table 2, where the average degree of under-explored nodes is much lower, making it easier for the models to predict links, resulting in a higher Recall.
+Additionally, in the under-explored partition, the rapid decline in Precision as K increases suggests that link prediction models are capable of discovering novel links that have not been actually explored by researchers in the test period.
+This highlights great opportunities for using “AI” to facilitate the expansion of the connections between AI methods and scientific challenges.
+Method # of novel (unseen) links
 @1 @3 @5 @10
-Train Node2vec 254 820 1,369 2,722
-LLM+RAG 870 2,502 3,983 6,775
-Train+Test Node2vec 4 396 899 2,203
-LLM+RAG 760 2,308 3,748 6,505
-Table 5. Number of novel links discovered by models compared with the training set and the whole dataset. The test data introduced 554 new links compared to the training data.
+Node2vec 8 132 432 1623
+LLM (Cluster) 243 867 1,523 3,209
+LLM (Cluster) 430 1,650 2,999 6,963
+LLM (Paper) 639 2,291 3,771 8,276
+Table 5.
+Number of novel links discovered by models compared with the training set (data up to 2022).
+The publications during 2023 and 2024 introduced 683 new links.
+For a fair comparison, in LLM (Cluster) we control the number of API calls to match that of LLM (Paper).
 4.4. Discovering Novel Links
-Previous discussions highlight the capability of link prediction models to uncover novel links.
-Building on this, we further investigate the newly discovered connections in response to Implication (I2).
-Compared to the training data up to 2022, human researchers identified 554 new links between scientific problems and AI methods during 2023 and 2024.
-In contrast, link prediction methods have demonstrated the potential to discover even more connections, especially as K increases, as shown in Table 5.
-Remarkably, the LLM+RAG approach demonstrates strong potential in generating new links, even with a relatively small number of predictions per data point (e.g., @1).
+Inspired by the capability of link prediction models to uncover novel links, we further explore the newly discovered connections in response to Implication (I2).
+Compared to the training data up to 2022, publications during 2023 and 2024 actually identified 683 new links between scientific problems and AI methods.
+In contrast, link prediction methods exhibit the potential to discover even more connections, especially as K increases, as shown in Table 5.
+Remarkably, the LLM-based approaches exhibit strong potential in generating new links, even with a relatively small number of predictions per data point (e.g., @1).
 This observation aligns with recent research suggesting that LLMs are capable of proposing novel research ideas (Si et al., 2024).
-Many of the predicted links involve applying Machine Learning, Deep Learning, and various types of Neural Networks (e.g., GNN, RNN, GAN), as well as Reinforcement Learning, to address a wide range of scientific challenges.
-In comparison, conventional link prediction methods like node2vec, which leverage cluster-level bipartite graph structures, exhibit high accuracy in identifying new links in the test data but show lower novelty compared to LLM+RAG.
-For instance, in the @1 predictions, the node2vec model successfully identifies 250 of the 254 new links (98.4%) that were also discovered by human researchers.
-The four remaining links, not identified by human researchers, include: Cardiac Pathophysiology – Matrix Factorization, Mycobacterium Tuberculosis – Contrastive Learning, Plant Stress Responses – Topological Data Analysis, and Prostate Cancer – Ensemble Learning.
-Some of these links have since been developed into publications (Wang et al., 2019; Tse et al., 2020; Palande et al., 2023; Patel et al., 2024), highlighting the model’s ability to suggest valuable connections ahead of human exploration.
+Interestingly, LLM (Paper) demonstrates a stronger capability in discovering novel links compared to LLM (Cluster), even when the number of API calls is controlled.
+This may be due to the greater diversity of paper-level inputs, leading to more varied outputs.
+Many of the predicted links have already materialized into publications, such as applying ‘Graph Neural Networks’ to ‘Protein Design and Function Prediction’ (Meller et al., 2023), using algorithms for ‘Molecular Dynamics Simulations’ to tackle challenges in ‘Polymer Dynamics and Chirality’ (Kuenneth and Ramprasad, 2023) and ‘Earth’s Mantle Dynamics’ (Li et al., 2023), or leveraging ‘Multimodal Learning’ for ‘Visual Perception and Neuroscience’ (Choi et al., 2023).
+Link prediction also highlights numerous unexplored but promising directions, such as using ‘Reinforcement Learning’ for ‘Carbon Emissions Mitigation’, applying ‘Graph Neural Networks’ to predict ‘RNA-Protein Interactions’, and using ‘Explainable AI’ to address challenges in ‘Artificial Intelligence Ethics‘.
+5. Related Work
+5.1. Qualitative Literature Reviews of AI4Science
+To identify potential applications of AI techniques in science, researchers typically turn to qualitative literature reviews of relevant papers.
+These reviews include general explorations of AI’s usage across various scientific domains (Wang et al., 2023; Leontidis, 2024), as well as more focused examinations in specific fields such as drug discovery (Dara et al., 2022; Mak et al., 2023; Blanco-Gonzalez et al., 2023; Qureshi et al., 2023), materials design (Sha et al., 2020; Li et al., 2020; Guo et al., 2021), and social sciences (Xu et al., 2024).
+While these reviews provide valuable insights, they often rely on interdisciplinary experts to manually identify and categorize scientific problems and potential AI methods, limiting their scope.
+Such an approach often could not offer a comprehensive and data-driven overview of the broader AI4Science landscape.
+5.2. Quantitative Literature Analysis
+In addition to heuristic and qualitative literature reviews, a few recent papers from the Science of Science community have focused on quantitatively analyzing the AI4Science publications.
+For instance, Gao and Wang (2023) discovered that AI use is prevalent across various scientific disciplines, and papers incorporating AI tend to receive a citation advantage.
+Duede et al.
+(2024) explored AI’s engagement in different fields, uncovering an “oil-and-water” phenomenon – while AI-engaged work is spreading across disciplines, it does not integrate well with other work within each specific field.
+Although these studies provide valuable insights into AI4Science, they primarily focus on areas where AI is already applied to the scientific problems, overlooking the broader and underexplored parts of the landscape.
+Moreover, these approaches often rely on pre-defined keywords or n-grams to tag AI4Science research, leading to a taxonomy that provides only a coarse representation of the AI and science space.
+This method fails to capture the actual semantics and nuances of the language, ultimately limiting the depth of analysis.
+By contrast, our study provides a more comprehensive view by analyzing literature from both AI and scientific venues over the past decade, encompassing not only AI4Science but also non-AI4Science publications.
+We leverage LLMs to extract detailed textual descriptions, enabling a nuanced and scalable data-driven categorization that facilitates deeper analysis.
+Our depiction of the landscape, quantitative analysis of discrepancies, and link prediction experiments offer fresh insights into the challenges and opportunities in bridging AI and science.
 6. Limitations, Challenges, and Future Work
 Potential bias in data selection.
-Our analysis focuses on publications from leading science and AI venues, specifically top-tier journals and conferences.
-While this ensures high-quality and influential work, it introduces a selection bias that may overlook significant contributions from less prominent venues, potentially resulting in an incomplete representation of the AI4Science landscape.
-To mitigate this bias, future work could expand the dataset to include a broader range of publications.
-Given that our analysis framework is flexible and can easily accommodate new data sources, such an expansion would be a valuable direction for achieving a more comprehensive understanding of the AI4Science field.
-Abstract-based extraction.
-The extraction of scientific problems, AI methods, and AI usages in this study rely solely on paper titles and abstracts, due to limited access to full texts and cost considerations.
-This approach aligns with the existing AI4Science literature analysis efforts, which also focus on the title and abstract as in the metadata (Gao and Wang, 2023; Duede et al., 2024).
-While this method provides useful insights, it may miss important details present in the full text of publications.
-For example, when extracting the AI methods used in the AlphaFold paper (Jumper et al., 2021), as illustrated in Fig.
-1, GPT identifies only the general use of a Neural Networks approach from the abstract.
-However, the specific use of the backbone Transformer model is detailed in the full text.
-In future work, efforts can be directed toward incorporating full-text analysis to enrich the extracted content and capture more nuanced AI methodologies used in scientific research.
-The definition of “AI4Science”.
-There is no universally accepted standard for defining the boundaries of the broad terms “AI”, “science”, and “AI4Science”, which presents a challenge for accurate classification.
-Previous studies have relied on keywords and human annotations (Gao and Wang, 2023; Duede et al., 2024), but this method can overlook less-known techniques and lacks the ability to capture the full semantic space of AI methods used in science.
-In contrast, our approach leverages the knowledge and capabilities of LLMs to classify AI4Science work, with human verification on a small-scale subset.
-To foster more consistent and scalable analyses, the AI4Science community should work toward developing standardized classification guidelines for automatic large-scale literature analysis.
+As an emerging and interdisciplinary field, cutting-edge and influential AI4Science studies often originate from prominent venues of general interest.
+For instance, groundbreaking work such as the AlphaFold (Jumper et al., 2021) was published in Nature.
+Based on this observation, we have focused on top-tier venues from the last decade.
+While this ensures high-quality and impactful contributions, it could introduce a selection bias that may overlook significant contributions from less prominent venues, potentially resulting in an incomplete representation of the AI4Science landscape.
+Given the flexibility of our analysis framework, researchers can integrate additional data sources from less-prominent venues to achieve a more representative view.
+LLM-based extraction from titles and abstracts.
+The LLM-based extraction in this study relies solely on paper titles and abstracts, due to limited access to full texts, challenge of processing PDF formats, relatively low signal-to-noise ratio in full texts, and cost considerations.
+Our approach is comparable to the existing AI4Science literature analysis efforts, which also focus on the title and abstract as in the metadata (Gao and Wang, 2023; Duede et al., 2024).
+While key research problems and methods are often described in abstracts, and human evaluation confirms the accuracy of LLM extraction (Appendix A.2), we may miss details presented in the full text of publications.
+Future work can build upon our general and flexible framework to incorporate full-text analysis, enriching the extracted content and capturing more nuanced AI methodologies used in scientific research.
 Evaluation of link prediction.
@@
 Limited exploration in link prediction models.
-Our primary goal in this study is not to optimize the link prediction model, but rather to use link prediction as a lens to explore the potential of connecting AI methods and scientific problems.
-As a result, we only included a few link prediction models, leaving room for further research to improve the performance of AI4Science link prediction or building recommender systems for AI methods or scientific problems.
+Our goal is not to optimize the link prediction model, but rather to use link prediction as a lens to demonstrate the potential of connecting AI methods and scientific problems.
+As a result, we only included a few straightforward link prediction models, leaving room for further research to improve the performance of AI4Science link prediction or building recommender systems for AI methods or scientific problems.
 7. Conclusion
-This paper aims to identify and bridge the gaps between AI and science in the context of AI4Science research.
-We introduce a comprehensive, large-scale dataset of AI4Science publications, with scientific problems and AI methods extracted using large language models (LLMs).
-Through quantitative analysis of this dataset, we uncover several key disparities: (1) Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods under-investigated; (2) The connectivity between scientific problems and AI methods is highly imbalanced, with certain nodes acting as “hubs” while other peripheral nodes are less connected; (3) The science and AI communities take distinct approaches to integrating AI into scientific research, prioritizing different problems and methods.
+This paper aims to depict the AI4Science landscape, identify current gaps between AI and science, while exploring potential ways to bridge them.
+We introduce a comprehensive, large-scale dataset of curated AI4Science publications, with scientific problems and AI methods extracted using large language models.
+Through quantitative analysis of this dataset, we uncover several key disparities: (1) Different AI and science subdomains exhibit varying degrees of engagement in AI4Science research, leaving a substantial number of scientific problems and AI methods underexplored; (2) The connectivity between scientific problems and AI methods is highly imbalanced, with certain nodes acting as “hubs” while other peripheral nodes are less connected; (3) The science and AI communities take distinct approaches to integrating AI into scientific research, prioritizing different problems and methods.
 We further investigate the potential and challenges of fostering AI4Science collaboration through the lens of link prediction.
@@
 We anticipate that our findings and tools will provide valuable insights to enhance interdisciplinary collaboration and accelerate scientific discovery through the integration of AI.
-Acknowledgements.
-The authors would like to extend their sincere appreciation to Shixuan Liu and Sicheng Shen for their contributions as annotators.
 Appendix A AI4Science Dataset
@@
     Table 6. Science Domain dataset statistics
-  • For AI domain, we obtained our data from MASSW (Zhang et al., 2024), which encompasses the following six top AI conferences: AAAI, IJCAI, ICLR, ICML, NeurIPS, WWW.
+  • For AI domain, we obtained our data from MASSW (Zhang et al., 2024), which encompasses the following seven top AI conferences: AAAI, IJCAI, ICLR, ICML, NeurIPS, WWW, SIGKDD.
 Only papers with both title and abstract available and within 2014 to 2024 are kept.
@@
     Conference #Pub. #AI4Science Pub.
-    NeurIPS 11,250 889
     AAAI 8,994 915
-    ICML 5,641 530
-    WWW 3,777 241
     IJCAI 2,953 271
     ICLR 2,907 205
-    Total 35,522 3,051
+    ICML 5,641 530
+    NeurIPS 11,250 889
+    WWW 3,777 241
+    SIGKDD 3,251 488
+    Total 38,773 3,539
     Table 7. AI Domain dataset statistics
@@
 Moreover, this shows that our LLM extraction results are more consistent compared with a group of human annotators.
+The high extraction accuracy indicated by our human evaluation also aligns with human evaluation results from recent studies (Zhang et al., 2024), where LLMs have been used to extract scientific workflows from titles and abstracts.
 Extraction acc. Human agreement F1
@@
 Embedding
-After obtaining the six key aspects of the papers, we first leverage the InstructionEmbedding model(instructor large (Su et al., 2022)) to obtain the semantic embeddings of the scientific problems and AI methods extracted.
+After obtaining the six key aspects of the papers, we first leverage the InstructionEmbedding model (instructor large)(Su et al., 2022) to obtain the semantic embeddings of the scientific problems and AI methods extracted.
+InstructorEmbedding is an advanced embedding model that generates task- and domain-specific text embeddings without additional training.
+Given our need to embed both scientific problems and AI methods, we use InstructorEmbedding with instructions below to provide contextualized embeddings tailored to our task rather than context-free embedding methods
 For scientific problems, we embed the problem keywords/keyphrases concatenated with the problem definition with the following instructions:
@@
 After we obtained the semantic embeddings, we used LargeVis (Tang et al., 2016) for dimension reduction followed by HDBSCAN (Campello et al., 2013) for clustering.
-Summarization
+HDBSCAN is a widely used density-based clustering algorithm designed for practical data mining tasks.
+As an extension of DBSCAN, it employs hierarchical clustering to identify clusters of varying densities and is more robust to parameter selection.
+A density-based clustering method allows us to identify clusters at appropriate granularities and avoids a top-down decision to predefine the number of clusters.
+Its efficiency ($O(n\log n)$ complexity) also makes it suitable for our dataset size.
+Labeling
 To have more consistent and generalizable knowledge of the clusters, we asked GPT (gpt-4o-2024-08-06) to summarize each cluster.
@@
 Then we provide the top words as well as examples of paper key aspects from a certain cluster to the LLM for summarization.
+Using LLMs with the below prompts to summarize clusters encourage the cluster labels to have the desirable properties (Mei et al., 2007): (1) understandable, (2) semantically relevant, (3) well covering the whole topic, and (4) distinguishable from other topics.
 For scientific problem clusters we use the following prompt:
@@
 Given a cluster of scientific problems/challenges, please help summarizing the cluster into a keyword or keyphrase less than 3 words.
-The top appearing words in the cluster together with some examples of extracted paper from that cluster will be provided.
+Moreover, top words from several nearby clusters are provided and your summary should accurately capture the essence of the cluster while differentiating it from neighboring clusters.
 Please summarize based on the provided information.
@@
 { top words }
-## Examples of scientific papers
-Below are some specific examples of scientific challenges in this cluster. For each sample the problem keyword, definition and discipline will be provided.
+## Examples of Scientific Problems
+Below are some specific examples of scientific problems in this cluster. For each sample the problem keyword, definition and discipline will be provided.
 { examples }
+## Top words from nearby clusters’ texual information
+{ Nearby clusters’ top words }
 ## Summarization Requirement
@@
 * If the summarization is not possible due to lack of information, output "N/A" when applicable.
+* The summarization should be specific enough to differentiate it from nearby clusters
+* Also provide summarization for nearby clusters
 ## Response Format
-Please output the summarization as a list, which has exactly *one* elements. The output should be in the format of a python list.
-["Keyword"]
+Please output the summarization as a list, which has exactly *(#nearby clusters+1) outputs. The output should be in the format of a python list.
+["Keyword", "Keyword neighbor 1",…]
 For AI method clusters we use the following prompt:
@@
 Given a cluster of Artificial Intelligence methods, please help summarizing the cluster into a keyword or keyphrase less than 3 words.
-The top appearing words in the cluster together with some examples of Artificial Intelligence metho from that cluster will be provided.
+The top appearing words in the cluster together with some examples of Artificial Intelligence methods from that cluster will be provided.
+Moreover, top words from several nearby clusters are provided and your summary should accurately capture the essence of the cluster while differentiating it from neighboring clusters..
 Please summarize based on the provided information.
@@
 ## Examples of Artificial Intelligence methods
-Below are some specific examples of Artificial Intelligence methods in this cluster. For each sample the problem keyword, definition and discipline will be provided.
+Below are some specific examples of Artificial Intelligence methods in this cluster. For each sample the method keyword and definition will be provided.
 {examples}
+## Top words from nearby clusters’ texual information
+{ Nearby clusters’ top words }
 ## Summarization Requirement
@@
 * If the summarization is not possible due to lack of information, output "N/A" when applicable.
+* The summarization should be specific enough to differentiate it from nearby clusters
+* Also provide summarization for nearby clusters
 ## Response Format
-Please output the summarization as a list, which has exactly *one* elements. The output should be in the format of a python list.
-["Keyword"]
-Merging Clusters
-To refine previous clustering results and obtain more consistent clusters, We perform cluster merging based on LLM summarization.
-If LLM gives the same summary for a group of adjacent clusters on the 2D projection map, we merge them into a large cluster with the same summary.
-After merging, we have $369$ scientific problems clusters and $345$ AI methods clusters as the final clusters for further analysis.
+Please output the summarization as a list, which has exactly *(#nearby clusters)* outputs. The output should be in the format of a python list.
+["Keyword", "Keyword neighbor 1",…]
 Appendix B Discrepancy Analysis
+Fig.
+5-6 illustrates the distribution of AI4Science and non-AI4Science publications across different communities, highlighting how AI and scientific fields engage with interdisciplinary research distinctively.
+Figure 5. Distribution of scientific problems of all publications, annotated with gpt-4o-2024-08-06 summary.
+Figure 6. Distribution of AI methods of all publications, annotated with gpt-4o-2024-08-06 summary.
+Well- and under-explored regions
+Table 12 list some top clusters of well- and under-explored scientific problems as well as AI methods.
+Scientific problems AI Methods
+Well-explored Under-explored Well-explored Under-explored
+Advanced Imaging Challenges Asymmetric Synthesis Deep Learning Models Attention Mechanisms
+Visual Perception and Neuroscience Magnetic Materials and Spintronics Convolutional Neural Networks Gradient Descent Methods
+Decision Making Forests and Climate Change Genomic Analysis Gradient-Based Methods
+Neural Network Challenges Viral Mechanisms Neural Networks Policy Gradient Methods
+Urban Traffic Management Evolutionary Paleontology Classification Methods Graph Embedding
+Alzheimer’s Disease Mechanisms Nanostructure Synthesis Computational Modeling Regularization Methods
+Statistical Inference Diabetes and Insulin Regulation Molecular Dynamics Simulations Transformer Architectures
+Depression and Anxiety Bacterial Infections and Immune Response Predictive Modeling Contrastive Learning
+Language Processing Cancer Signaling Pathways Advanced Imaging Techniques Graph Representation
+Genetic Variants and Traits Cancer Drug Resistance Computational Drug Screening Adaptive Algorithms
+Table 12. Scientific problems and AI methods that are well- and under-explored by AI4Science research. Cluster names are ranked by the number of total publications. Only top 10 clusters are shown.
 Bipartite graph node degrees.
 Table 13 presents the nodes with the highest degrees in the bipartite graph, representing the most connected scientific challenges and AI methods in the AI4Science landscape.
-These high-degree nodes act as key “hubs,” indicating their central role in bridging AI techniques with scientific problems
+These high-degree nodes act as key “hubs,” indicating their central role in bridging AI techniques with scientific problems.
 Discrepancies between the AI and science communities.
-Fig.
-4-5 illustrates the distribution of AI4Science and non-AI4Science publications across different communities, highlighting how AI and scientific fields engage with interdisciplinary research distinctively.
-Table.
-Figure 4. Distribution of scientific problems of all publications, annotated with gpt-4o-2024-08-06 summary.
-Figure 5. Distribution of AI methods of all publications, annotated with gpt-4o-2024-08-06 summary.
 (a) Distribution of scientific problems of scientific journal papers, annotated with gpt-4o-2024-08-06 summary.
@@
 Scientific Problem Node Deg. AI Method Node Deg.
-1 Neural Network Challenges 229 Machine Learning 218
-2 Statistical Inference 162 Deep Learning 149
-3 Language and Auditory Processing 149 Genomic Analysis 122
-4 Optimization Problems 135 Advanced Imaging and Spectroscopy 91
-5 Social Media Dynamics 134 Computational Biology 79
-6 Decision-Making Processes 126 Protein Design 76
-7 Market Economics 118 Neural Networks 70
-8 Game Theory 115 Molecular Modeling 65
-9 Dynamical Systems 115 Image Processing 64
-10 Imaging Challenges 106 Classification Methods 60
-11 Causal Inference 105 Data Analysis 58
-12 Visual and Sensory Perception 97 Bayesian Inference 57
-13 Robotic Actuators 92 Materials Informatics 55
-14 Urban and Sustainability Challenges 84 Unsupervised Learning 52
-15 Optical and Wireless Communication 79 Wearable Sensor Technology 51
-16 Complex Network Dynamics 78 Predictive Modeling 51
-17 Scientific Policy Challenges 76 Drug Screening 49
-18 Neural Imaging 72 Deep Neural Networks 49
-19 Traffic Prediction 69 Convolutional Neural Networks 46
-20 Protein Structure and Design 63 Artificial Intelligence Techniques 44
+1 Neural Network Challenges 164 Machine Learning 140
+2 Urban Traffic Management 122 Genomic Analysis Methods 137
+3 Electronic Health Data Challenges 102 Advanced Imaging Techniques 98
+4 Imaging Challenges 89 Computational Drug Screening 97
+5 Statistical Inference 89 Molecular Dynamics Simulations 88
+6 Decision Making 86 Protein Structure Design 88
+7 Language Processing 85 Deep Learning Models 84
+8 Dynamical Systems 84 Neural Networks 62
+9 Visual Perception and Neuroscience 82 Classification Methods 62
+10 Market Pricing and Allocation 77 Molecular Structural Design 61
+11 Neural Imaging Challenges 74 Deep Neural Networks 57
+12 Single-cell RNA Sequencing 71 Predictive Modeling 53
+13 Soft Robotics 69 Convolutional Neural Networks 52
+14 Drug Discovery 67 Computational Modeling 52
+15 Protein Design and Function Prediction 67 Unsupervised Learning 50
+16 Social Media Dynamics 67 Image Segmentation 50
+17 Mechanical Properties of Materials 64 Data Analysis 41
+18 Causal Inference 62 Deep Learning Framework 39
+19 Optimization Problems 60 Data-driven Approaches 39
+20 Game Theory Challenges 58 Recurrent Neural Networks 37
 Table 13. Nodes in the bipartite graph (Fig. 4a) with the highest degrees.
@@
 In the science community In the AI community In the science community In the AI community
-1 Neural Network Challenges Neural Network Challenges Machine Learning Deep Learning
-2 Imaging Challenges Social Media Dynamics Deep Learning Machine Learning
-3 Language and Auditory Processing Statistical Inference Genomic Analysis Data Analysis
-4 Visual and Sensory Perception Language and Auditory Processing Advanced Imaging and Spectroscopy Neural Networks
-5 Protein Structure and Design Game Theory Protein Design Game Theory
-6 Materials Design Market Economics Neural Networks Bayesian Inference
-7 Artificial Intelligence Ethics Optimization Problems Computational Biology Causal Inference
-8 Statistical Inference Decision-Making Processes Molecular Modeling Probabilistic Modeling
-9 Single-Cell RNA Sequencing Dynamical Systems Optical Computing Reinforcement Learning
-10 Single-Cell Omics Causal Inference Materials Informatics Optimization Methods
-11 Decision-Making Processes Robotic Actuators Image Processing Predictive Modeling
-12 Cancer Detection Imaging Challenges Neuroimaging Techniques Sequential Decision Making
-13 Gene Expression Regulation Optical and Wireless Communication Memory-augmented Systems Convolutional Neural Networks
-14 Genetic Variants Traffic Prediction Drug Screening Spatio-temporal Analysis
-15 Dynamical Systems Urban and Sustainability Challenges Wearable Sensor Technology Fair Resource Allocation
-16 Chemical Reaction Dynamics Fair Resource Allocation Unsupervised Learning Regret Minimization
-17 Protein-Protein Interactions Voting and Elections Molecular Dynamics Policy Optimization
-18 Wearable Tactile Sensors Visual and Sensory Perception Classification Methods Auction Mechanism Design
-19 Antibiotic Resistance Scientific Policy Challenges Deep Neural Networks Classification Methods
-20 Motor Control and Learning Complex Network Dynamics Bayesian Inference Generative Models
+1 Neural Network Challenges Urban Traffic Management Machine Learning Deep Learning Models
+2 Imaging Challenges Neural Network Challenges Genomic Analysis Methods Machine Learning
+3 Single-cell RNA Sequencing Electronic Health Data Challenges Deep Learning Models Reinforcement Learning
+4 Mechanical Properties of Materials Language Processing Molecular Dynamics Simulations Attention Mechanisms
+5 Visual Perception and Neuroscience Statistical Inference Advanced Imaging Techniques Convolutional Neural Networks
+6 Protein Design and Function Prediction Market Pricing and Allocation Computational Drug Screening Neural Networks
+7 Cancer Detection Soft Robotics Protein Structure Design Recurrent Neural Networks
+8 Genetic Variants and Traits Dynamical Systems Molecular Structural Design Graph Neural Networks
+9 Decision Making Decision Making Brain Imaging Techniques Policy Gradient Methods
+10 Language Processing Game Theory Challenges Computational Modeling Multitask Learning
+11 Gene Regulatory Networks Causal Inference Optical AI Methods Classification Methods
+12 Drug Discovery Drug Discovery Deep Neural Networks Transformer Architectures
+13 Cancer Mutation Studies Optimization Problems Wireless Sensor Technology Spatio-Temporal Analysis
+14 Mitochondrial DNA Mutations Social Media Dynamics Neural Networks Deep Reinforcement Learning
+15 Protein-Protein Interactions Imaging Challenges Image Segmentation Predictive Modeling
+16 Neural Imaging Challenges Urban Dynamics Unsupervised Learning Causal Inference
+17 Cryo-electron microscopy Neural Imaging Challenges Classification Methods Self-Supervised Learning
+18 Tactile Sensing Technology Visual Perception and Neuroscience Convolutional Neural Networks Graph Convolutional Networks
+19 Chemical Reaction Dynamics Sparse Signal Processing Predictive Modeling Unsupervised Learning
+20 Chromatin Organization Energy Management Neuromorphic Devices Latent Variable Models
 Table 14.
@@
 Publication year range 2014-2022 2023-2024
-Number of All publications 138,577 20,718
-AI4Science publications 5,829 1,195
-Pub. in well-investigated sci. clusters 3,503 512
-Pub. in under-investigated sci. clusters 1,129 379
-Pub. in well-investigated AI clusters 3,038 687
-Pub. in under-investigated AI clusters 1,336 186
-Avg. deg. of Scientific problem nodes 14.4 3.7
-Well-investigated sci. nodes 59.5 8.9
-Under-investigated sci. nodes 5.3 2.5
-AI method nodes 14.7 4.7
-Well-investigated AI nodes 25.6 7.4
-Under-investigated AI nodes 7.8 1.8
+Number of All publications 141,639 21,017
+AI4Science publications 6,287 1,225
+Pub. in well-explored sci. clusters 4,262 594
+Pub. in under-explored sci. clusters 1,180 392
+Pub. in well-explored AI clusters 3,076 742
+Pub. in under-explored AI clusters 1,387 198
+Avg. deg. of Scientific problem nodes 11.3 2.8
+Well-explored sci. nodes 37.7 6.9
+Under-explored sci. nodes 3.9 1.4
+AI method nodes 12.4 3.0
+Well-explored AI nodes 16.2 4.9
+Under-explored AI nodes 7.2 1.3
 Table 15. The training and test splits of data.
@@
 We examine two approaches to link prediction based on different data representations:
-  • Bipartite graph link prediction (“$P_{i}\to M_{i}$” or “$M_{i}\to P_{i}$”): In this approach, scientific problems and AI methods are represented as high-level cluster labels, serving as nodes in a bipartite graph (as defined in Sec.
-LABEL:).
+  • Cluster-level link prediction (“$P_{i}\to M_{i}$” or “$M_{i}\to P_{i}$”): In this approach, scientific problems and AI methods are represented as high-level cluster labels, serving as nodes in a bipartite graph.
 This captures more generalized connections between similar scientific problems and AI methods.
-  • Generative link prediction (“$p_{i}\to m_{i},u_{i}$” or “$m_{i}\to p_{i},u_{i}$”): Here, scientific problems and AI methods are expressed in detailed text descriptions rather than coarse cluster labels.
+  • Paper-level link prediction (“$p_{i}\to m_{i},u_{i}$” or “$m_{i}\to p_{i},u_{i}$”): Here, scientific problems and AI methods are expressed in detailed text descriptions rather than coarse cluster labels.
 This approach leverages textual generative models, such as large language models (LLMs), to capture more fine-grained distinctions and predict links based on textual data.
@@
 Evaluation metrics.
-For bipartite link prediction, we evaluate performance using Recall, Precision, and F1 scores @K, where K represents the number of generated candidate links, as is standard in the literature (Al Hasan et al., 2006; Yang et al., 2015).
+For cluster-level link prediction, we evaluate performance using Recall, Precision, and F1 scores @K, where K represents the number of generated candidate links, as is standard in the literature (Al Hasan et al., 2006; Yang et al., 2015).
 Particularly, for a source node $s$, supposing in the test data, it is connected to a set of target nodes $T(s)$, and the model generates $K$ predictions $P(s)$.
@@
 (3) Recall :=|$:=\frac{|\{t\in P(s)\mid t\in T(s)\}|}{|T(s)|}.$
-For link generation, we adopt three widely used evaluation metrics for text generation: ROUGE (Lin, 2004), BLEURT (Sellam et al., 2020), and the cosine similarity of text embeddings.
+For paper-level link prediction, we adopt three widely used evaluation metrics for text generation: ROUGE (Lin, 2004), BLEURT (Sellam et al., 2020), and the cosine similarity of text embeddings.
 The best scores @K are reported for each of these metrics.
@@
 $cos(r,e)=\frac{<r,e>}{||r||\ ||e||}$
-LLM generation setup
-For generative link prediction, we consider both Sci $\to$ AI and AI $\to$ Sci generations:
+LLM (Paper) link prediction setup
+For paper-level generative link prediction, we consider both Sci $\to$ AI and AI $\to$ Sci generations:
   • Sci $\to$ AI. Given the scientific problem description, generate the AI method that can solve the scientific problem.
@@
 Prompts
-For generative link prediction with LLMs (LLM+RAG), we use the following prompt [^1] For brevity only Sci $\to$ AI RAG prompt is provided:
+For paper-level link prediction with LLMs, we use the following prompt [^1] For brevity only Sci $\to$ AI RAG prompt is provided:
 ## Background and Task Description
@@
 "AI Usage": "…"
-For graph-level link prediction with LLMs (LLM+Graph), we use the following prompt:  [^2] For brevity only Sci $\to$ AI RAG prompt is provided:
+For cluster-level link prediction with LLMs (LLM (Cluster)), we use the following prompt:  [^2] For brevity only Sci $\to$ AI RAG prompt is provided:
 ## Background and Task Description
@@
 C.2. Experiment Results
-Tables 15-19 summarize the detailed LLM link generation experiment results with different models and evaluation metrics. These results show that:
+Tables 17-21 summarize the detailed LLM link generation experiment results with different models and evaluation metrics. These results show that:
   • Including RAG examples helps the model make more informed predictions, as all evaluation metric results improve when we increase the examples given from $n=1$ to $n=5$.
   • While Imitation can achieve higher rouge scores, it has less semantic similarity compared with LLM generations.
+  • Table 16 show that the performance of LLM (Paper) remains consistent across different models used, suggesting a low risk of data leakage.
+However, the performance of LLM (Cluster) declines, likely due to GPT-3.5’s limited ability of reasoning and processing long contexts.
+Setting Model Precision Recall F1
+@1 @3 @5 @10 @1 @3 @5 @10 @1 @3 @5 @10
+Sci $\to$ AI LLM (Paper,gpt-3.5-turbo-0125) 0.273 0.191 0.175 0.141 0.052 0.081 0.109 0.134 0.087 0.114 0.134 0.138
+LLM (Paper,gpt-4o-2024-08-06) 0.282 0.227 0.200 0.156 0.058 0.106 0.129 0.155 0.096 0.145 0.157 0.155
+LLM (Cluster,gpt-3.5-turbo-0125) 0.134 0.168 0.167 0.137 0.014 0.065 0.108 0.180 0.025 0.094 0.131 0.155
+LLM (Cluster,gpt-4o-2024-08-06) 0.352 0.300 0.238 0.159 0.053 0.101 0.134 0.176 0.093 0.151 0.171 0.167
+AI $\to$ Sci LLM (Paper,gpt-3.5-turbo-0125) 0.516 0.421 0.401 0.210 0.099 0.127 0.148 0.148 0.166 0.195 0.216 0.174
+LLM (Paper,gpt-4o-2024-08-06) 0.377 0.323 0.307 0.296 0.065 0.104 0.127 0.163 0.111 0.158 0.180 0.211
+LLM (Cluster,gpt-3.5-turbo-0125) 0.136 0.082 0.120 0.102 0.009 0.023 0.053 0.088 0.017 0.036 0.074 0.095
+LLM (Cluster,gpt-4o-2024-08-06) 0.201 0.168 0.143 0.129 0.016 0.042 0.053 0.092 0.030 0.068 0.077 0.107
+Table 16. Link prediction results comparison between gpt-3.5-turbo-012 and gpt-4o-2024-08-06.
 Sci $\to$ AI ROUGE-1-F BLEURT CosineSim
@@
 Sci $\to$ AI All test data 0.290 0.297 0.305 0.314 0.417 0.432 0.440 0.444 0.887 0.894 0.897 0.900
-Well-investigated 0.284 0.300 0.308 0.317 0.418 0.439 0.446 0.447 0.882 0.890 0.893 0.885
-Under-investigated 0.301 0.316 0.324 0.333 0.429 0.451 0.459 0.460 0.896 0.904 0.906 0.902
+Well-explored 0.284 0.300 0.308 0.317 0.418 0.439 0.446 0.447 0.882 0.890 0.893 0.885
+Under-explored 0.301 0.316 0.324 0.333 0.429 0.451 0.459 0.460 0.896 0.904 0.906 0.902
 AI $\to$ Sci All test data 0.255 0.267 0.266 0.276 0.387 0.395 0.402 0.407 0.818 0.829 0.838 0.845
-Well-investigated 0.243 0.251 0.258 0.262 0.383 0.389 0.395 0.399 0.827 0.840 0.848 0.852
-Under-investigated 0.246 0.255 0.265 0.264 0.390 0.398 0.406 0.410 0.822 0.833 0.844 0.848
+Well-explored 0.243 0.251 0.258 0.262 0.383 0.389 0.395 0.399 0.827 0.840 0.848 0.852
+Under-explored 0.246 0.255 0.265 0.264 0.390 0.398 0.406 0.410 0.822 0.833 0.844 0.848
 Table 21. LLM+RAG performances on text generation metrics.
+Appendix D Supplementary Materials
+We include our dataset, code, analyses, and experiment results in the anonymous repository(https://github.com/charles-pyj/Bridging-AI-and-Science) as supplementary materials to this paper, which include:
+  • The AI4Science dataset with publication extractions and cluster labels.
+  • Tables of cluster statistics for generating Fig. 3.
+  • Link data of the bipartite graph and new links introduced by LLM-based models.
+  • Degree information of scientific problem nodes and AI method nodes.
+  • Human annotations and evaluation results with two human annotators.
+  • Code implementations for visualizations, link predictions, metrics, and data analysis, which generates the main results of the paper.